### v1

Abstract
3D direct numerical simulation of electrolysis is applied to investigate the growth and detachment of bubbles at electrodes. The moving gas-liquid interface is modeled employing the VOF-based method. To ensure the accuracy of the simulations, a mesh-independence study has been performed. The simulations include the growth phase of the bubbles followed by their detachment from the electrode surface and the results are validated with analytical models and experimental data. The bubble growth is diffusion controlled leading to the scaling $R=2\beta t^{1/2}$, but the growth exponent is overpredicted by our simulation during initial stage. Furthermore, it is proved that the nucleation sites of the bubble strongly influence gas transport by measuring the relevant Sherwood number. Finally, we investigate the effects of contact angle and nucleation sites on bubble detachment behavior, and compare the detachment radius with Fritz’s formula, the results show a good agreement, confirming that buoyancy is the dominant driving force. As the nucleation sites increase, the induced bubble coalescence accelerates the bubble detachment. Taken together, these findings give us valuable insights into improving gas bubble removal and enhancing overall electrolysis efficiency.
keywords:
[…]
1 Introduction
The production of green hydrogen through water electrolysis is expected to be an important technology in achieving global ”Net-zero emission” targets (Turner, 2004; Holladay et al., 2009; Dawood et al., 2020). However, it is slow and inefficient in many situations. The attached bubbles in such electrochemical devices reduce the efficiency of electrolyzer systems by blocking the active electrode sites or by increasing the ohmic resistance (Swiegers et al., 2021). Maintaining a bubble-free electrode surface is therefore crucial for high efficient $\mathrm{H_{2}}$ production. Consequently, a detailed understanding of bubble evolution dynamics is highly desirable for developing strategies to enhance water electrolysis efficiency. Analytical solutions for this problem are not generally available, and only a few exact solutions can be derived using extremely simplified assumptions (Epstein and Plesset, 1950; Scriven, 1959). However, they can still serve as valuable references for experimental and numerical studies (Glas and Westwater, 1964; Brandon and Kelsall, 1985; Dapkus and Sides, 1986; van der Linde et al., 2017; Taqieddin et al., 2017; Soto et al., 2018; Angulo et al., 2020; Zhang and Lohse, 2023).
Several experiments have been done, and they confirmed that bubble growth is mainly driven by mass transport due to the gradient of the dissolved hydrogen concentration. Glas and Westwater (1964) performed a simple experiment for hydrogen bubbles generated on a flat electrode. The experimental findings demonstrate that the asymptotic growth of electrochemically generated bubbles follows the same functional relationship given by the analytical solution for a suspended bubble growing in a supersaturated liquid, see Scriven (1959). Recent studies have already investigated this problem from different scales. Macroscopic processes are mainly convection phenomena in the electrolyte solution, which is observed at the electrolytic cell scale. Li et al. (2018) visualized experimentally on two-phase flow at the anode side of a proton exchange membrane electrolyzer. They understood that the inlet velocity does not affect bubble growth when the temperature and current density are constant. Moreover, they investigated gas bubble (hydrogen and oxygen) dynamics in a single-channel electrolyzer (Li et al., 2019). There, they showed that the bubble detachment diameter is inversely proportional to the flow velocity. Because of the strong influence of electrolyzer geometry on macro scale bubbly flow, Hreiz et al. (2015) studied several different electrochemical configuration sets. In the following work (Abdelouahed et al., 2014), Laser Doppler Velocimetry (LDV) measurements were conducted to observe bubble curtain behavior. Other experimental studies revealed that the bubble curtain velocity is a function of the average electric current (Hine and Murakami, 1980), the mass transfer coefficient is varied for horizontal and vertical electrodes (Fouad and Sedahmed, 1972), and bubble coverage is also a function of current density (Vogt and Balzer, 2005). Vogt (2012) has mentioned that when the bubble coverage tends to unity, the reaction is totally blocked. For the micro-scale, the microphysical phenomena surrounding individual bubbles constitute the overall behavior of the electrolysis process, and consequently, understanding the micro-scale bubble hydrodynamic is vital to reveal the governing mechanisms behind electrolysis. As the studied scale becomes smaller, however, the difficulty of the experiment rises dramatically. Several studies confirm that flow convection in the micro-area is caused by bubble growth, detachment, and coalescence (micro-convection) strongly influences the hydrogen diffusion boundary layer and changes mass transfer (Stephan and Vogt, 1979; Vogt and Balzer, 2005). van der Linde et al. (2018) used a specially designed electrode with pillars and pits to control micro-bubble formation, measuring the bubble radius and attachment force to verify the Fritz radius. Apart from bubble formation and growth, Bashkatov et al. (2024) investigated two modes of coalescence-induced bubble detachment using a dual platinum microelectrode system. By combining high-speed imaging and electrochemical analysis, they demonstrated the importance of bubble-bubble interactions in the departure process. The impact of micro-convection induced by bubble dynamics on mass transfer is significant. Burdyny et al. (2017) demonstrated that micro-convection enhances mass transport primarily by reducing the departure diameter of bubbles from the electrode surface. With advanced visualization techniques, the gas concentration released in the liquid by bubbles can be investigated (Dani et al., 2007; Francois et al., 2011). However, experimental measurements are generally expensive and limited by the available measuring techniques, which usually provide global quantities (e.g., global mass transfer rate, rising velocities), and do not give information about local details, such as small scale multiphase flow field and local mass transfer rate.
Due to the presence of large amounts of gas, conventional optical techniques are unable to detect many critical aspects of multiphase flow fields. Computational Fluid Dynamics (CFD) has become an important tool for the comprehensive study of electrolytic complex multiphase flows (Hawkes et al., 2009; El-Askary et al., 2015). The significant increase in computational power has made direct numerical simulation an important alternative method for studying the detailed dynamics of mass transfer between fixed or deformable interfaces.
The individual bubble growth, multiple bubbles coalescence, and detachment that occur in this region have been numerically studied for a long period. It is confirmed that the diffusion-driven bubble growth dominates the growth stage (Soto, 2019). The bubble detachment radius is approximated by the Fritz radius (Vogt et al., 2004). It can be derived by the force balance acting on the growing and detaching bubbles. Besides, the mass transfer behind the phenomenon still needs to be discussed. van der Linde et al. (2017); van der Linde et al. (2018) simulated the hydrogen concentration field around a growing hydrogen bubble in acidic electrolysis, using a body-fitted axisymmetric finite difference method. Sepahi et al. (2022) employed an Immersed Boundary Method (IBM) to simulate the mass transport for bubbles in gas-evolving electrolysis, indicating that the net transport within the system is governed by the effective buoyancy driving induced by the rising bubble. Very few simulations of deformable-bubble growth, detachment, and rise in the context of electrolysis have been performed using modern sharp-interface methods such as VOF, level set, or front tracking. Moreover, some authors used VOF methods without mass transfer (Lafmejani et al., 2017). Most previous studies, such as those by Sepahi et al. (2022) and Khalighi et al. (2023), assumed undeformable spherical bubbles, neglecting bubble deformation that may arise due to lateral flow during detachment. Simulations of deformable bubbles without the spherical bubble assumption have not been performed very often due to the capillary time-step constraint. Thus, a rigid sphere approximation saves considerable CPU time. Although these models do not apply to the design of real systems, they still provide helpful information and reference solutions for the validation of other models. Recently, Gennari et al. (2022) developed a VOF-based phase-change model for diffusion-driven mass transfer problems using the one-fluid method and a novel algorithm to extrapolate the discontinuous velocity field across the interface to improve interface advection accuracy. This method is applied to study the growth of deformable bubbles on planar electrodes.
In this paper, we investigate all these effects using direct numerical simulation (DNS), which means that we fully resolve all relevant scales of the hydrodynamic and concentration boundary layers. We take advantage of the free code repository Basilisk (http://basilisk.fr/).
2 Configuration and numerical methods
To simplify the complicated mathematical description of a bubble growing at an electrode in the presence of a surrounding flow, consider a two-phase gas-liquid system represented in fig. 1. We assume a constant room temperature, dilute liquid solutions ($\mathrm{KOH}$, 0.5 mol/L ), and no evaporation of water. Besides, in order to avoid further complications, self-ionisation of water is disregarded due to its low equilibrium constant at room temperature. The equation for the electric potential remains relatively simple. At the cathode, the reaction process is as follows,
$\mathrm{2H_{2}O}+\mathrm{2e^{-}}\rightarrow\mathrm{2H_{2}+2OH^{-}}.$ (1)
In the present work, only the reaction at the cathode is considered.
Figure 1: Schematic representation of the two-phase electrochemical system with relevant chemical reactions and boundary conditions at the cathode.
2.1 Problem set-up
The bubble growth and detachment are investigated first under axisymmetry. Due to the geometric axisymmetry, an axisymmetric simulation can also effectively address the problem when studying the growth of a single bubble on a circular electrode in addition to a full 3D simulation. A sketch of the axisymmetric setup and mesh grid is shown in fig. 2. The axisymmetric condition is applied at the left boundary (Z axis), while an outflow boundary condition is set on the top boundary; the other boundaries are treated as no-slip wall. The computational domain is a square. The domain’s characteristic length is $L_{0}=25L_{ref}$, where the reference length is the initial bubble diameter $D_{b}$. The bubble is initialized in simulations with the diameter $D_{b}=0.0127$ mm, The electrode is the flat end of a wire with a diameter of $D_{e}=10D_{b}=0.127$ mm, oriented along the axis of symmetry (Z axis).
As for the more complex configuration, "mutiple nucleation sites", the geomertic axisymmetry does not exist anymore. To address the mutual interaction, a 3D configuration is needed. The multiple nucleation sites are equally spaced at the electrode center, as illustrated in fig. 3.
The initial liquid is set to be saturated, and the saturation ratio is $\zeta=c_{0}/c_{sat}=1$, where $c_{sat}$ is the hydrogen concentration in the saturated liquid and $c_{0}$ is the initial concentration of dissolved gas near the electrode surface. As mentioned in section 1, the production of dissolved gas at the electrode walls creates a locally supersaturated region ($\zeta>1$). It drives the growth of bubbles formed from microscopic pits on the electrode surface due to the heterogeneous nucleation, see Jones et al. (1999); van der Linde et al. (2017). According to the experiment, there should be no bubble present at initial time $t=0$. However, using the VOF method requires the volume fraction of gas need to be initialized. We wait for a nucleation time before computing the volume change. The bubble size is fixed during this stage ($t<t_{n}$), while the concentration of dissolved hydrogen continues to increase. This approach enables the development of a concentration field around the bubble by the time nucleation occurs, better reflecting experimental observation. The nucleation time varies due to many factors, like the electrode material and the current density. In present work, we set the nucleation time $t_{n}=0.02$ s, which is observed from Glas and Westwater (1964). The control parameters for the electrolytically generated bubbly flow are the cathodic current density, and the contact angles for different wettability of the electrode surface. The current density $I$ can give the molar flux of hydrogen ($\mathrm{H_{2}}$) by Faraday’s law,
$J=\frac{I}{2F},$ (2)
where $I=i/A$, $i$ is the total electric current, $A$ is the cross-section area ($A=\pi/4D_{e}^{2}$); $F$ is Faraday’s constant ($F=96485.3$ As/mol). To account for the flux of H2 across the active area of the electrode, a Neumann boundary condition for the gas concentration is applied to the electrode wall (r-axis for axis-symmetric simulation, XY plane for 3D simulation),
$\frac{\partial c}{\partial z}$ $=\frac{J}{D}\qquad\text{for }r<\frac{D_{e}}{2},$ (3)
$\frac{\partial c}{\partial z}$ $=0\qquad\text{for }r>\frac{D_{e}}{2}.$ (4)
The boundary influx from the bubble-free region directly leads to an increase of the local concentration of dissolved hydrogen, leading to the development of a hydrogen concentration layer. This process facilitates hydrogen bubble formation, driving its growth and detachment. Simulations performed in the present work are classified into two groups. The configurations are listed in Table 1. The axisymmetric configuration is mainly used for numerical method verification by comparing the bubble growth with analytical and experimental results. The 3D configuration is designed to study the bubble coalescence and detachment.
Figure 2: Sketch of the axisymmetric simulation setup.
Figure 3: Sketch of the 3D simulation setup
Configuration
Single bubble (Axisymmetric) $I(Am^{-2})$
No. $\theta\degree$ $z\times r(L_{ref})$ mesh level 100
1 90 $25\times 25$ 8 300
2 90 $25\times 25$ 9 1000
3 90 $25\times 25$ 10 -
4 35 $25\times 25$ 8 -
5 35 $25\times 25$ 9 -
6 35 $25\times 25$ 10 -
3D
No. $\theta\degree$ $x\times y\times z(L_{ref})$ nucleation sites -
7 90 $25\times 25\times 25$ 1 -
8 90 $25\times 25\times 25$ 2 -
9 90 $25\times 25\times 25$ 4 -
10 35 $25\times 25\times 25$ 1 -
11 35 $25\times 25\times 25$ 2 -
12 35 $25\times 25\times 25$ 4 -
Table 1: Simulation parameters for cases with a single bubble and mutiple bubble nucleation sites. All the simulation are performed at 3 different current densities, as listed in the last column.
The physical parameters remain the same for all the simulation cases (see Table 2). The only exception is that gravity acceleration is set to zero for the axisymmetric configuration. To reduce the computational cost due to the time step limitation of the surface tension scheme, $\sigma$ is decreased by a factor of $10^{-4}$. Besides, the density of hydrogen is relatively small, and the large density ratio for this system ($\rho_{c}/\rho_{d}=12450$) would be a problem for the numerical simulation. It is observed that a large density ratio will slow down the convergence of the multigrid solver. To circumvent this problem, the hydrogen density is increased to $\rho_{d}=0.8\,\text{kg/m}^{3}$, the molar mass in increased by the same factor. so the volume change is not influenced ($\Delta V\propto M/\rho_{d}$). The zero gravity settings are only applied for the axisymmetric configuration. When it comes to the investigation of bubble detachment, gravity acceleration is applied at the direction of - Z in 3D configuration with the value of $9.8\mathrm{m/s^{2}}$.
Symbol Properties Value Unit
$\rho_{c}$ Electrolyte density 996 [$\mathrm{kg/m^{3}}$]
$\rho_{d}$ Hydrogen density 0.8 [$\mathrm{kg/m^{3}}$]
$g$ Gravity acceleration 9.8 [$\mathrm{m/s^{2}}$]
$M$ Molar mass of hydrogen $2.0\times 10^{-2}$ [$\mathrm{kg/mol}$]
$c_{0}$ H2 initial concentration 0.002 [$\mathrm{mol/m^{3}}$]
$c_{\text{sat}}$ H2 saturated concentration 0.02 [$\mathrm{mol/m^{3}}$]
$\mu_{c}$ Electrolyte viscosity $8.32\times 10^{-4}$ [$\mathrm{kg/(m\cdot s)}$]
$\mu_{d}$ Hydrogen viscosity $8.96\times 10^{-6}$ [$\mathrm{kg/(m\cdot s)}$]
$\nu_{c}$ Electrolyte kinematic viscosity $8.35\times 10^{-7}$ [$\mathrm{m^{2}/s}$]
$\sigma$ Surface tension $7.5\times 10^{-2}$ [$\mathrm{N/m}$]
$F$ Faraday’s constant $96485.3$ [$\mathrm{C/mol}$]
$D$ Hydrogen diffusion coefficient $7.38\times 10^{-9}$ [$\mathrm{m^{2}/s}$]
$L_{\text{ref}}$ Reference length $1.27\times 10^{-5}$ [$\mathrm{m}$]
$D_{b}$ Initial bubble diameter $1.27\times 10^{-5}$ [$\mathrm{m}$]
$D_{e}$ Electrode diameter $1.27\times 10^{-4}$ [$\mathrm{m}$]
Table 2: Physical properties in SI units.
2.2 Non-dimensional numbers
The most basic non-dimensional number related to the transport of $\mathrm{H_{2}}$ is the Sherwood number. The production rate of hydrogen flux is constant at the electrode, and $J$ is a constant value in time. The transport of hydrogen results in a surface-averaged concentration, $\bar{c_{e}}$ at the electrode surface. To illustrate the transport, we compare the $\bar{c_{e}}$ with the concentration value in the bulk liquid, for which we adopt the top of the domain. Since the electrolyte solution is initially saturated ($\zeta=1$), the initial value $c_{0}=c_{sat}$, it yields the Sherwood number of production of hydrogen,
$\overline{\mathrm{Sh_{e}}}=\frac{JL_{ref}}{D(\bar{c_{e}}-c_{sat})}.$ (5)
The bar symbol is used to mark the surface-averaged response parameters. By introducing the boundary layer thickness $\delta_{\mathrm{H_{2}}}=D\Delta c/J$, this Sherwood number could be expressed as $\mathrm{Sh}=D_{b}/\delta_{\mathrm{H_{2}}}$. For pure diffusion, the $\delta_{\mathrm{H_{2}}}$ ultimately reaches the electrolysis cell height irrespective of the current density so that the Sherwood could maintain stability for all the cases without the effect of the bubbles.
For the mass transfer considering local volume change, the hydrogen transport to the bubble will be considered. This leads to the expression for the bubble Sherwood number,
[…]
where $A_{\Sigma}$ is the interface surface area, $4\pi R$, and the concentration difference between the electrode and bubble interface, $c_{e}-c_{\Sigma}$. According to eq. 31, the concentration of the interface $c_{\Sigma}=c_{sat}$. Since we assume the configuration with an ambient temperature and pressure, the mass transfer rate across the interface can be defined as:
[…]
where $\mathcal{R},T_{0},P_{0}$ are the universal gas constant, ambient temperature, and pressure, respectively. Then it leads to a rather simple expression of bubble Sherwood number,
$\overline{\mathrm{Sh_{e}}}=\frac{P_{0}}{\mathcal{R}T_{0}}\frac{2R}{D(\bar{c_{e}}-c_{\Sigma})}\frac{dR}{dt}$ (8)
Apart from this crucial Sherwood number, which measures the mass transfer, other nondimensional numbers are also vital as control parameters for physics. First, the Schmidt number compares momentum diffusion to mass diffusion,
$\mathrm{Sc}=\frac{\nu_{c}}{D},$ (9)
Where the $\nu_{c}$ is the kinematic viscosity of the electrolyte. For these configurations with gravity added, the dynamics of bubbles in a gravitational field can be described by the Galilei and Bond numbers. The Galilei number compares the gravitational and viscous forces and is defined as:
$\mathrm{Ga}=\sqrt{\frac{\rho_{c}gD_{b}^{2}}{\nu_{c}^{2}}}.$ (10)
The ratio of gravitational and capillary forces is estimated by the Bond number,
$\mathrm{Bo}=\frac{\rho_{c}gD_{b}^{2}}{\sigma}.$ (11)
2.3 Governing equation
A model of the microscopic process involves the solution of the two-phase incompressible Navier-Stokes equations with phase change, surface tension, gravity, and contact line dynamics on the wall. Due to the constant temperature assumption, no equation for the thermal energy is needed.
The gas phase is called $\Omega_{d}(t)$, for the disperse phase, and the liquid phase is called $\Omega_{c}(t)$, for the continous phase. These two subdomains are separated by an infinitely thin interface $\Sigma(t)$. The entire domain is given by $\Omega=\Omega_{d}(t)\cup\Omega_{c}(t)\cup\Sigma(t)$. The normal vector $n_{\Sigma}$ at the interface points into $\Omega_{d}(t)$.
For each phase, the governing equation of an incompressible flow system in the absence of mass transfer reads,
$\nabla\cdot\mathbf{u}$ $=0\quad\text{in }\Omega\backslash\Sigma,$ (12)
$\begin{split}\partial_{t}(\rho\mathbf{u})+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u})&=-\nabla p+\nabla\cdot(2\mu\mathbf{D})\\ &\quad+\rho\mathbf{g}\quad\text{in }\Omega\backslash\Sigma.\end{split}$ (13)
where the density $\rho$ and viscosity $\mu$ remain constant in $\Omega_{d}$ and $\Omega_{c}$. The eq. 12 is the continuity equation, where $\mathbf{u}$ represents the velocity field. In the balance of momentum eq. 13, $p$ is the static pressure, $\mathbf{D}$ is the deformation tensor, and $g$ represents body force, which is gravity acceleration in this system. So, $a$ is replaced by $g$ for the following sections. The eq. 12 and eq. 13 are valid everywhere in the domain except at the interface, where additional conditions are needed (Tryggvason et al., 2011). The continuity equation requires that the amount of mass that leaves one phase $\Omega_{d}(\Omega_{c})$ must be transferred to another phase $\Omega_{c}(\Omega_{d})$ since the infinitely thin interface region can not store any mass. It results in a jump condition across the interface,
$\rho(\mathbf{u}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}=0,$ (14)
where the jump notation has been introduced; $\mathbf{u_{\Sigma}}$ is the interface velocity and $\dot{m}$ is the mass transfer rate $\mathrm{kg/m^{2}s}$. The second jump condition is derived by applying the conservation of momentum to a control volume with an infinitely small thickness around the interface, and it reads,
$\mathbf{u}\otimes(\mathbf{u}-\mathbf{u_{\Sigma}})+p\mathbf{I}-2\mu\mathbf{D}\cdot\mathbf{n_{\Sigma}}=\sigma k\mathbf{n_{\Sigma}}+\nabla_{\Sigma}\sigma.$ (15)
where $\mathbf{I}$ is the unit tensor, $\sigma$ is the surface tension and $k$ is the curvature of the interface. In the problem of hydrogen bubble growth, the interface is considered to have uniform surface tension, besides a no-slip boundary condition is applied at the interface, eq. 15 can be further simplified, see Fleckenstein and Bothe (2015).
$\left(p\mathbf{I}-2\mu\mathbf{D}\|\cdot\mathbf{n_{\Sigma}}\right)=\sigma k\mathbf{n_{\Sigma}}$ (16)
The numerical method used in this work for interface transport is the Volume of Fluid (VOF) method, and is combined with one-fluid formulation of the governing equations. In the one-fluid approach, the jump condition eq. 14 and eq. 16 are replaced by source terms that act at the interface as singularities ($\delta$ function), a single set of Navier-Stokes equations for the entire domain $\Omega$ is solved.
$\nabla\cdot\mathbf{u}$ $=\dot{m}\left(\frac{1}{\rho_{d}}-\frac{1}{\rho_{c}}\right)\delta_{\Sigma},$ (17)
$\partial_{t}\mathbf{u}+\nabla\cdot(\mathbf{u}\otimes\mathbf{u})$ $=\frac{1}{\rho}\left[-\nabla p+\nabla\cdot(2\mu\mathbf{D})\right]+\frac{\sigma k\mathbf{n_{\Sigma}}}{\rho}\delta_{\Sigma},$ (18)
where the $\delta_{\Sigma}$ is the surface Dirac function which has a nonzero value only at the interface. So the system of eq. 17 eq. 18 is valid in the whole domain. Then, to determine the location of the interface, a marker function is required. The Heaviside function serves this purpose,
$H(x,t)=\begin{cases}1,&\text{if }x\in\Omega_{c},\\ 0,&\text{if }x\in\Omega_{d}.\end{cases}$ (19)
The transport equation for Heaviside function $H(x,t)$ can be obtained from the following integral balance for a control volume $V$.
[…]
where the second term on the LHS represents the convective transport, and the last term is a source term that accounts for the mass transfer across the interface, which is null when $\dot{m}=0$. Converting the surface integral to volume integral, we can write it into differential form,
$\partial_{t}H+\nabla\cdot(H\mathbf{u})+\frac{\dot{m}}{\rho_{c}}\delta_{\Sigma}=0.$ (21)
eq. 21 is then served to compute the volume fraction, see section 2.5.
2.4 Concentration transport equation
This model can be used to compute the concentration field of the soluble hydrogen in a two-phase system with mass transfer by applying the two-scalar method of Fleckenstein and Bothe (2015). In the present study, we focus on pure incompressible gas bubbles, and we assume that no electrolyte species exists in the disperse phase (i.e., the electrolyte is not volatile). The system contains two different species: hydrogen (denoted with subscript 1) and electrolyte liquid (denoted with subscript 2). The overall mass transfer is entirely given by the transfer of the hydrogen species. It is worth pointing out that we do not need to solve the mass balance in the dispersed phase since no mixture exists inside the bubbles (pure hydrogen gas). The mass balance of hydrogen in the continous side $\Omega_{c}$ reads
$\partial_{t}\rho^{1}+\nabla\cdot(\rho^{1}\mathbf{u^{1}})=R,$ (22)
where $\rho^{1}$ is the partial density of hydrogen, $\mathbf{u^{1}}$ denotes the hydrogen velocity. $R$ is the reaction term that can model chemical reactions. It is coupled with the jump condition for the conservation of mass,
$\|\rho^{1}(\mathbf{u^{1}}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}||=||\dot{m}^{1}||=0.$ (23)
The average phase density and velocity of are derived from the respective species terms,
$\rho_{c}=\rho^{1}+\rho^{2},$ (24)
and
$\rho_{c}\mathbf{u}=\rho^{1}\mathbf{u^{1}}+\rho^{2}\mathbf{u^{2}},$ (25)
where the superscript $2$ denotes the species electrolyte. Then the transport equation of hydrogen for incompressible flow can be written as
$\partial_{t}\rho^{1}+\mathbf{u}\cdot\nabla\rho^{1}+\nabla\cdot J^{1}=R,$ (26)
and the diffusive flux of hydrogen is
$J^{1}=\rho^{1}(\mathbf{u^{1}}-\mathbf{u}).$ (27)
The mass transfer rate of hydrogen can be derived from eq. 23,eq. 14.
$\begin{split}\dot{m}^{1}&=\rho^{1}(\mathbf{u}-\mathbf{u}_{\Sigma})\cdot\mathbf{n}_{\Sigma}+\rho^{1}(\mathbf{u}^{1}-\mathbf{u})\cdot\mathbf{n}_{\Sigma}\\ &=\frac{\rho^{1}}{\rho_{c}}\cdot\dot{m}+J^{1}\cdot\mathbf{n}_{\Sigma}\end{split}$ (28)
which shows that the mass transfer contains both a convective term and a diffusive term. Under the assumption of dilute liquid solutions, the diffusive flux can be well modeled by Fick’s law of diffusion
$J^{1}=-D^{1}\nabla\rho^{1},$ (29)
where the $D^{1}$ is the hydrogen diffusion coefficient. Combining the eq. 29 and eq. 28, the mass transfer rate of hydrogen reads
$\dot{m}=-\frac{M^{1}D^{1}}{1-\frac{\rho^{1}}{\rho_{c}}}\frac{\partial c^{1}}{\partial\mathbf{n_{\Sigma}}},$ (30)
where the molar concentration has been introduced, i.e. $c^{1}=\rho^{1}/M^{1}$ and $M^{1}$ is the molar mass. It gives the mass transfer rate evaluated from the continuous side of the interface, where the concentration of the dissolved species is generally variable. Since we have only hydrogen, which can be transferred between phases, in the following, we will omit the species indicator as we will always refer to the concentration of the soluble hydrogen in the liquid phase $\Omega_{c}$.
To compute the concentration gradient, we need the hydrogen concentration at the liquid side of the interface. For a gas-liquid system at equilibrium, we can employ Henry’s law to compute the concentration on the liquid side of the interface.
$(c_{c})_{\Sigma}=\frac{(c_{d})_{\Sigma}}{\mathrm{He}},$ (31)
where $\mathrm{He}$ is the Henry coefficient, and it is taken as a constant for the present work. So, the hydrogen concentration at the liquid side of the interface $(c_{c})_{\Sigma}$ is immediately computed. $(c_{d})_{\Sigma}=\rho_{d}/M$ is a constant as the density should be equal everywhere inside the bubble ($\Omega_{d}$).
For numerical integration, we rewrite the mass balance eq. 26 for the hydrogen diffusion process. As we discussed, the mass balance will be done in the continuous region $\Omega_{c}$ and for a non-reactive flow ($R=0$). Since the hydrogen dissolved in $\Omega_{c}$ is the only species that we need for mass transportation. The phase indicator will be omitted in the following, i.e., $c=c_{c}$.
[…]
Combine the jump condition eq. 23, the final differential form could be deduced.
$\partial_{t}c+\mathbf{u}\cdot\nabla c=\nabla\cdot\left(D\nabla c\right)-\frac{\dot{m}}{M}\delta_{\Sigma}.$ (33)
2.5 Numerical methodology
The phase change model is derived from the work of Gennari et al. (2022). The governing equation shown in section 2.3 is solved by open source solver Basilisk (HTTP://basilisk.fr/) which is a finite-volume P.D.E. solver on adaptive Cartesian grids. Using adaptive mesh refinement (AMR) in regions with large gradients makes the approach particulally suitable for multiscale processes such as interfacial flows. In interfacial flows, a fine mesh is typically required around the gas-liquid interface but not the entire domain. The shape of the domain is always a square $L_{0}\times L_{0}$ in axisymmetric configuration (cube $L_{0}\times L_{0}\times L_{0}$ in 3D). The grid is organized following a hierarchical quadtree (octree) structure, where each cell can be further divided into four child cells (eight in 3D), and a level is assigned to each cell according to the relative tree structure. The root cell is at level 0, and its size $\Delta$ is the same as the whole numerical domain ($\Delta=L_{0}$). A generic cell at level $l$ has a resolution of $\Delta(l)=\frac{L_{0}}{2^{l}}$. The grid structure in Basilisk allows neighboring cells to vary by up to one level, meaning each cell edge/face can communicate with no more than two finer edges/faces. An example of a quadtree grid with different levels of refinement and the relative tree structure is shown in fig. 4.
Figure 4: An example of quadtree discretisation (left) together with its logical representation (right). The level of the cells in the tree is also given (Popinet, 2012).
The VOF method is one of the most widely used numerical approaches for the modeling of two-phase immiscible fluids. The starting point for the derivation of the VOF approach is the one-fluid formulation presented in section 2.3, which is vital to the phase change problem. The volume fraction of the continuous phase is defined as
[…]
and the value of $f_{c}$ is within the set $[0,1]$, depending on the amount of liquid in the cell.
$f_{c}=\begin{cases}0&\text{if the cell is pure gas},\\ 1&\text{if the cell is pure liquid},\\ [0,1]&\text{if the cell is mixed}.\end{cases}$ (35)
The volume fraction of the dispersed phase $f_{d}$ is implicitly described by the relationship $f_{c}+f_{d}=1$. So, only one transport equation of H needs to be solved. By applying the incompressibility constraint, the integrated form of eq. 21 is
[…]
According to the finding of Scardovelli and Zaleski (1999), the integration of the transport eq. 36 is performed in two steps, namely the reconstruction step and the propagation step. First, the interface is approximated with a line/plane in each interfacial cell. Second, the fluxes of volume fraction across the cell boundaries are computed, and eq. 36 is integrated in time. The geometric reconstruction of the interface is based on the piecewise linear interface construction (PLIC) method, where the interface is approximated as a line (plane) in axisymmetric configuration (3D).
3 Implementation
In this section, we show the simulation results of several different configurations; as described in section 2.1. This numerical study is designed to simulate the experiments conducted by Glas and Westwater (1964). The experimental result confirm the driven force of bubble growth is diffusion and measure the bubble growth rate. The comparsion between following numerical and experimental growth rate will verify the accuracy of simulation. The physical properties corresponding to an alkaline solution (typically used in industrial water electrolysis) are shown in Table 2.
The simulation results are presented with dimensional units for comparison with experimental measurements. According to the analytical and experimental work, the growth of electrochemically generated bubbles follows the same functional relationship as the solution for bubble growth in a supersaturated liquid described by the Scriven model. Different stages or regimes for bubble growth have been characterized by a power law: growth controlled by inertia ($R\propto t$), by diffusion ($R\propto t^{1/2}$) or by reaction limitation ($R\propto t^{1/3}$) , see (Angulo et al., 2020). The growth behavior ($R\propto t^{1/2}$) reflects a standard analytical solution for diffusive bubble growth. To validate the specific relationship, we perform a series of axisymmetric simulations first.
3.1 Axisymmetric simulation validation
In this section, we assume there is only one bubble (with an initial diameter $D_{b}$) on the cathode. In fig. 5, the left-hand side of the snapshots illustrates hydrogen evolution during the bubble’s growth stage, where color represents the dissolved hydrogen concentration in the liquid phase. For visualization consistency, the gas phase inside the bubble will also be colored, but will not be calculated (based on the assumption of constant pressure inside the bubble) and will remain at 1. The same map is applied to the rest of the snapshots in the following. Hydrogen is produced at the electrode surface, so hydrogen saturation near the electrode wall is more significant than that in the bulk liquid. Completing the visualization, the right side of the snapshots displays the symmetric grid cell and interface, clearly demonstrating the mesh refinement near the gas-liquid interface.
The growth rate of a single bubble confirms the result of the diffusion-controlled power law ($R\propto t^{1/2}$). As shown in fig. 6, a mesh independence study demonstrates that mesh refinement level 8 will be enough for interface tracking. A logarithmic plot of the bubble growth for different current densities between $t=0.02\mathrm{s}$ and $t=0.2\mathrm{s}$ of our simulation (mesh refinement level 8) is shown in fig. 7.
In our simulation case, the hydrogen flux is applied before the bubble nucleation time ($t_{n}=20\mathrm{ms}$), with the average hydrogen concentration on the electrode wall increasing in time. After nucleation, the concentration difference drives bubble growth through the diffusion process. As a result, the bubble growth behavior shows some discrepancies from the analytical solution during the early stages of diffusion. For smaller current density $I=100\mathrm{A/m^{2}}$, the growth exponent is smaller than $1/2$. For much larger current density $I=1000\mathrm{A/m^{2}}$, the growth follows the power law ($R\propto t^{0.8}$) better. In a similar simulation using the immersed boundary method, see Khalighi et al. (2023), the growth exponent 0.8 is also observed in single bubble growth.
Scriven’s result is valid for a spherically symmetric concentration field, bubble radius much larger than the initial radius, constant solubility, and, far away from the bubble, constant concentration and zero velocity. This solution is confirmed in cases where the thickness $\delta_{\mathrm{H_{2}}}$ of the diffusion boundary layer, surrounding the bubble is small compared with the diameter of the bubbles.
Given the constraints of the analytical solution, the discrepancies in our simulation can be explained. The hydrogen flux at the boundary determined by current density influences the growth exponent increasing the averaged hydrogen concentration over the electrode wall. At the beginning stage, the growth exponent is not close to $1/2$. The bubble radius is relatively small compared to the thickness of the diffusion boundary layer. For high current densities $I=1000\mathrm{A/m^{2}}$, the enhanced hydrogen flux elevates the supersaturation level and creates a strong diffusion boundary layer. The diffusion layer continuously provides $\mathrm{H_{2}}$ for bubble growth, which by consequence leads a much faster $\mathrm{H_{2}}$ production than consumption and a growth exponent exceeding $1/2$. Conversely, lower current densities reduce hydrogen availability, can not fulfill the consumption of bubble growth, leading to sub $1/2$ exponents.
After the initial growth stage, as the bubble radius increases, the $\delta_{\mathrm{H_{2}}}$ is much smaller compared with the bubble radius. The growth exponents for different boundary condition are all approaching $1/2$, which indicates that the overall growth is diffusion controlled.
Figure 5: The dissolved hydrogen concentration for a bubble growing at different times (The current density is $I=\mathrm{1000A/m^{2}}$).The color lengend represents the numerical $\mathrm{H_{2}}$ concentration.
Figure 6: Bubble radius growth for bubbles with $\theta=90\degree$ for different current densities. The units of $I$ are $\mathrm{A/m^{2}}$
Figure 7: The growth bubble radius slope and growth exponent variation.
To further compare the numerical results with experimental data, additional simulations are performed with a smaller contact angle ($\theta=35^{\circ}$), which is the same value from the experimental study(see Glas and Westwater (1964)). From the analytical solution $R=2\beta t^{1/2}$, the growth rate $\beta$ could be deduced. The results in terms of the influence of current density on the growth rate are reported in fig. 8. The first comparison indicates that the growth rate increases with rising current density, attributed to the high concentration of dissolved hydrogen near the electrode. This observation is supported by both the numerical simulations conducted by Gennari et al. (2022) and experimental results from Glas and Westwater (1964). The observation can be explained by the current density controlling the hydrogen flux at the electrode, which drives bubble growth.
Figure 8: Comparison of growth rate between experimental and numerical results.
We find that a bubble with a contact angle of $\theta=90^{\circ}$ grows more rapidly than one with a contact angle of $\theta=35^{\circ}$. This difference can be explained by considering the geometric difference of different contact angles. The mass transfer through diffusion occurs only through the interface between the gas and liquid phase. In the $\theta=90^{\circ}$ scenario, the bubble elongates more in the r-direction (parallel to the electrode wall), thereby increasing its exposure to the high-concentration region close to the electrode. So, the mass transfer for the bubble with $\theta=90\degree$ would be higher. A qualitative comparison between $\theta=90\degree$ and $\theta=35\degree$ is shown in fig. 9 and fig. 10.
It is notable that the simulation of a single bubble overpredicts the growth rate compared to the experimental measurements. This discrepancy is observed not only in the present simulation but also in the results from Gennari et al. (2022). In fact, there are at least four active nucleation sites on the electrode surface from the experiment of Glas and Westwater (1964). This suggests that experimental observations of single bubble growth are generally influenced by neighboring bubbles. Depending on the material and surface properties of the electrode, if multiple nucleation sites are active, the dissolved hydrogen in the liquid is consumed simultaneously by all of them. Consequently, the relative growth rates are smaller than in the case of a single nucleation site, where the entire amount of $\mathrm{H_{2}}$ produced by the electrode is available for a single bubble on the wall. This mechanism explains the overprediction in the simulations. To further investigate the effect of variations in nucleation site distribution, a 3D simulation would be necessary, as the scenario would no longer be axisymmetric.
Figure 9: Bubble shape and contour of $\mathrm{H_{2}}$ concentration for $\theta=90\degree$.
Figure 10: Bubble shape and contour of $\mathrm{H_{2}}$ concentration for $\theta=35\degree$.
3.2 3D simulation extension
To extend the numerical scheme to 3D simulation, we set the domain as a cubic (fig. 3). And the boundary condition would be modified to suit the domain. The electrode will be placed in the middle of the domain bottom(XY plane), which will give the same boundary condition of hydrogen flux as axisymmetric simulation, see eq. 4. The top plane(XY) will be set as an outflow boundary condition, and the surrounding area (XZ and YZ planes) adopt a no-slip wall boundary condition.
In this section, a single bubble is modeled in 3D to validate the 3D applicability of the numerical method. We compare the bubble growth prediction of the 3D simulation to that of the corresponding axisymmetric simulation. Simulations of configurations No.1 and No.7 (Table 1) with three different current densities are listed for comparison. The contact angle is $90\degree$. Same bubble growth slope from section 3.1 are observed in fig. 11. Since the convergence has been verified in section 3.1, the good agreement with the axisymmetric simulation results proves the compatibility of 3D simulation.
Figure 11: Bubble radius growth for bubbles with $\theta=90\degree$ for different current densities. The units of $I$ are $\mathrm{A/m^{2}}$.
4 Result and discussion
4.1 Multiple bubbles growth
There is a notable decrease in the growth rate and exponent of hydrogen bubbles when multiple nucleation sites are active from the experiment of Glas and Westwater (1964). To reproduce this key observation, this section numerically investigates the mutual bubble interactions for multiple nucleation sites. Instead of a single bubble, we model 2 bubbles and 4 bubbles growing at the electrode. To highlight the effect of nucleation sites, a higher boundary flux is applied, with the current density set to $\mathrm{1000A/m^{2}}$. The simulation results for the average bubble radius growth of two and four bubbles are presented in fig. 12. The growth of multiple nucleation sites is mutually suppressed in the stage before bubble coalescence. During this stage, hydrogen is consumed by all the bubbles simultaneously, leading to a reduced growth rate and exponent. The more active nucleation sites, the slower the growth, as shown in the comparison between experimental and simulation results in fig. 13. However, an unexpected behavior of the growth seems puzzling. Intuitively, merged bubbles should approximate the growth dynamics of a single nucleation site, leading to convergent growth across all cases. Contrary to this assumption, even after bubble coalescence($t>t_{m}$), differences in the growth persist.
Figure 12: Bubble growth of configurations No.7 to No.9. The current density is $I=\mathrm{1000A/m^{2}}$ and the slope represents the growth exponent.
Figure 13: Influence of nucleation sites on the growth rate.
A possible explanation lies in the relative position of bubbles within the hydrogen concentration layer. As shown in fig. 12, the trend remains consistent: the more active nucleation sites, the slower the growth, regardless of the stage (before or after bubble coalescence).
(a) Single nucleation site at merge time $t_{m}$.
(b) Four nucleation sites at merge time $t_{m}$.
Figure 14: Comparison of bubble growth between different nucleation sites. The current density is 1000 A/m2. The view direction is the same as in fig. 3. The hydrogen contour on the intersecting plane (YZ) in the middle of the bubble is transferred to the back plane.
The key reason for the different growth behavior after coalescence is the bubble radius. Observations indicate that the radius of the merged bubble in the four-nucleation-site case is larger than in the dual-site case, while the bubble in the single-site case has the smallest radius. Despite this, the hydrogen concentration layer remains nearly identical across different nucleation-site cases at the same time. This is because the hydrogen flux at the boundary is sufficiently high, making the amount of hydrogen consumed by bubble growth negligible. Then, the crucial factor is immersion depth: bigger bubbles immersed shallower into the hydrogen concentration layer, leading to a lower mass transfer rate, which means a smaller growth exponent, as illustrated in fig. 12.
Supporting evidence emerges from direct comparisons: at the same time $t_{m}$, the four-nucleation-site configuration produces marginally larger bubbles than the single-site case, see fig. 14. However, the constant hydrogen flux induces a countervailing effect as illustrated in the snapshots fig. 14, the bubble surface is colored by the mass transfer rate. A higher mass transfer rate slows the increase of the hydrogen concentration boundary layer for the four-nucleation-site case, enabling shallower bubble immersion after merge. This feedback mechanism drives a gradual decrease in growth exponent, ultimately causing convergence across nucleation-site configurations. As shown in fig. 12, the growth exponent gradually approaches a slope of 0.5.
To quantitatively assess mass transfer, we analyze the time-dependent bubble Sherwood number defined in eq. 8, which quantifies mass transport efficiency to the bubble. The initial condition $\zeta=1$, delay the plot until $t>t_{n}$, which is shown in fig. 15. It should be noticed that multiple nucleation sites exhibit elevated $\mathrm{Sh_{b}}$ initially. A decrease occurs after bubble merge($t_{m}$), which causes the $\mathrm{Sh_{b}}$ to be smaller than the one for the single nucleation site. The plot of $\mathrm{Sh_{b}}$ cross-verified the explanation of growth exponent above. First, the mutual suppression. While total mass transport increases with nucleation sites, individual bubble transport efficiency is inhibited through competitive $\mathrm{H_{2}}$ consumption. Second, the immersion depth: lower after-merge $\mathrm{Sh_{b}}$ directly correlates with diminished mass transfer rates for multiple nucleation sites, consistent with the immersion depth arguments.
Figure 15: Corresponding Sherwood number for simulations of configurations No.7 to No.9.
4.2 Bubble detachment
The removal of growing bubbles is crucial for enhancing the efficiency of the electrolysis process (Angulo et al., 2020). The bubble detachment radius is approximately equal to the Fritz radius (Stephan and Vogt, 1979)
$R_{det}=0.6\theta\sqrt{\frac{\sigma}{(\rho_{c}-\rho_{d})g}},$ (37)
where $\theta$ the contact angle in radians. The general contact angle for the bubble at the electrode would be smaller than $\pi/2$. This analytical solution is derived by balancing the buoyancy force for a perfect sphere-shaped bubble $F_{b}=(4/3)\pi R^{3}(\rho_{c}-\rho_{d})g$ with the capillary force for a bubble attached to the electrode. The contact angle controls the magnitude of surface tension effects and varies in connection with contact line dynamics, which are still, in many circumstances, an object of investigation. Molecular dynamics simulations with either Lennard-Jones intermolecular potentials or more realistic water molecule models give very different results, see Lācis et al. (2022). Therefore, in the present study, the theoretical value is adopted as a reference to estimate the order of magnitude rather than for precise comparison.
To begin with, a simulation of single-bubble detachment is shown in fig. 16. The numerical configuration refers to the No.1 and No.4 in Table 1. As observed, once the mass transfer is considered (at $t=20\mathrm{ms}$), the bubble starts to grow. Initially, the bubble remains nearly spherical due to its small size, which results in a negligible buoyancy force. However, as the bubble radius increases, deformation becomes evident, particularly after $t=52.1\mathrm{ms}$. Eventually, buoyancy overcomes surface tension, leading to detachment from the electrode surface.
Figure 16: Frames show the detachment evolution($\theta=90^{\circ}$).
As previously mentioned, eq. 37 establishes a relationship between the detachment radius and the contact angle. To further investigate this dependency, another case with a contact angle $\theta=35\degree$ was also simulated, see fig. 17. Compared to the case $\theta=90\degree$, the detachment time is much shorter, and the detachment radius is much smaller. fig. 18 presents a good agreement between the numerical simulation and the Fritz radius.
Figure 17: Frames show the detachment evolution($\theta=35^{\circ}$).
Figure 18: Bubble detachment radius versus contact angle for a single bubble
Additionally, it has been confirmed that mutual interactions between bubbles influence the detachment process. Bashkatov et al. (2024) reported that bubble coalescence in the presence of dual bubbles can lead to significantly earlier detachment and a smaller detachment radius compared to cases where only buoyancy effects are considered. To validate this finding, we perform a series of simulations involving bubble detachments for multiple nucleation sites. The results of 3D simulations with four nucleation sites are presented in fig. 19. Furthermore, varying the number of nucleation sites demonstrates that bubble coalescence indeed accelerates detachment, as shown in fig. 20. As a result, the bubble detachment radius is also smaller, see fig. 21.
A plausible mechanism behind this acceleration could be the following. The coalescence-induced shape perturbations generate an initial momentum favoring premature detachment from the electrode surface. It is observed in both our simulation and Bashkatov’s experiment that the coalescence of bubbles results in an initial jump-off of the merged bubble. This can be attributed to the released surface energy during coalescence (Zhang et al., 2024). The released energy is partly dissipated by the bubble oscillations, working against viscous drag. When in the proximity to the surface, the remaining energy is converted to kinetic energy, driving the resultant (merged) bubble to jump off the electrode (Lv et al., 2021). This kinematic energy can cause bubble departure at smaller radii than in the buoyancy-driven scenario, as observed in our simulation results. However, the capillary wave due to energy dissipation is not observed in our simulation because of the mesh resolution constraint. Further numerical investigations are needed to distinguish the respective roles of buoyancy and interfacial effects in this process.
Figure 19: snapshots of the detachment evolution for 4 nucleation sites (Configuration No.9).
Figure 20: Bubble detachment times for different nucleation sites and contact angles.
Figure 21: Bubble detachment radius for different nucleation sites and contact angles.
5 Conclusion
This study investigates the mechanisms of bubble growth, coalescence, and detachment at horizontal electrodes through three-dimensional direct numerical simulations. A convergence study is conducted to verify numerical accuracy, with results compared against experimental data. Key parametric analyses include the effects of bubble contact angle and multi-bubble interactions. The following conclusions are drawn:
Bubble growth dynamics The bubble contact angle significantly influences growth behavior by dictating the geometric position of the bubble within the evolving hydrogen concentration profile. This directly modulates local mass transport rates. For multi-bubble systems, mutual suppression of mass transport occurs due to simultaneous hydrogen consumption by neighboring bubbles, as quantified by reductions in Sherwood number (fig. 15). While simulations overpredict growth rates compared to experimental observations (fig. 13), they adhere closely to the scaling law of Scriven’s solution ($R\propto t^{1/2}$).
Bubble detachment and coalescence behavior Detachment radii align with Fritz’s formula, demonstrating a decrease for smaller contact angles. Multi-bubble systems exhibit earlier detachment with reduced radii compared to isolated bubbles. t is worth noting that coalescence reduces the detachment time, thus potentially improving the mass transfer efficiency of the electrochemical system.
Prospective The present work focuses on single growth-detachment cycles due to computational constraints. Tracking deformable interfaces remains resource-intensive, limiting simulations to short time spans ($t\sim 0.2$s). While comparable to prior studies using rigid bubble models (Sepahi et al., 2022), broader parametric investigations (e.g., current density variations, contact angle distributions) are needed to strengthen statistical conclusions. Future efforts are needed to develop a new scheme with subgrid model, which will help to enable long-term simulations.
Additionally, it should be noted that the experimental data used for comparison are not fully aligned with the study’s objectives. We will design and implement a more sophisticated experimental setup that will provide observational proof of our work (e.g., continuous single bubble detachment), thus increasing the reliability and impact of the 3D numerical study.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.
Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement number 883849). We thank the French national GENCI supercomputing agency and the relevant supercomputer centers for their grants of CPU time on massively parallel machines, and their teams for assistance and the use of Irene-Rome at TGCC.

### v4

@@
 Abstract
-3D direct numerical simulation of electrolysis is applied to investigate the growth and detachment of bubbles at electrodes.
+Three-dimensional direct numerical simulation of electrolysis is applied to investigate the growth and detachment of bubbles at electrodes.
 The moving gas-liquid interface is modeled employing the VOF-based method.
 To ensure the accuracy of the simulations, a mesh-independence study has been performed.
-The simulations include the growth phase of the bubbles followed by their detachment from the electrode surface and the results are validated with analytical models and experimental data.
-The bubble growth is diffusion controlled leading to the scaling $R=2\beta t^{1/2}$, but the growth exponent is overpredicted by our simulation during initial stage.
-Furthermore, it is proved that the nucleation sites of the bubble strongly influence gas transport by measuring the relevant Sherwood number.
-Finally, we investigate the effects of contact angle and nucleation sites on bubble detachment behavior, and compare the detachment radius with Fritz’s formula, the results show a good agreement, confirming that buoyancy is the dominant driving force.
-As the nucleation sites increase, the induced bubble coalescence accelerates the bubble detachment.
+The simulations include the growth phase of the bubbles, followed by their detachment from the electrode surface, and the results are validated with analytical models and experimental data.
+The bubble growth is diffusion-controlled, leading to the scaling $R\propto t^{1/2}$, but our simulation overpredicts the growth exponent during the initial stage.
+We further demonstrate that the number of nucleation sites significantly affects gas transport, as quantified by the Sherwood number.
+The influences of contact angle and nucleation site on bubble detachment are also examined.
+The predicted detachment radius varies linearly with contact angle, consistent with Fritz’s linear relation between the volume-equivalent radius and contact angle, confirming that the surface tension is the dominant attachment force.
+Finally, as the nucleation sites increase, the induced bubble coalescence accelerates the bubble detachment.
 Taken together, these findings give us valuable insights into improving gas bubble removal and enhancing overall electrolysis efficiency.
@@
 1 Introduction
-The production of green hydrogen through water electrolysis is expected to be an important technology in achieving global ”Net-zero emission” targets (Turner, 2004; Holladay et al., 2009; Dawood et al., 2020).
+The production of green hydrogen through water electrolysis is expected to be an important technology in achieving global net-zero emissions (Turner2004; Holladay2009; Dawood2020).
 However, it is slow and inefficient in many situations.
-The attached bubbles in such electrochemical devices reduce the efficiency of electrolyzer systems by blocking the active electrode sites or by increasing the ohmic resistance (Swiegers et al., 2021).
-Maintaining a bubble-free electrode surface is therefore crucial for high efficient $\mathrm{H_{2}}$ production.
+The attached bubbles in such electrochemical devices reduce the efficiency of electrolyzer systems by blocking the active electrode sites or by increasing the ohmic resistance (Swiegers2021).
+Maintaining a bubble-free electrode surface is therefore crucial for highly efficient $\mathrm{H_{2}}$ production.
 Consequently, a detailed understanding of bubble evolution dynamics is highly desirable for developing strategies to enhance water electrolysis efficiency.
-Analytical solutions for this problem are not generally available, and only a few exact solutions can be derived using extremely simplified assumptions (Epstein and Plesset, 1950; Scriven, 1959).
-However, they can still serve as valuable references for experimental and numerical studies (Glas and Westwater, 1964; Brandon and Kelsall, 1985; Dapkus and Sides, 1986; van der Linde et al., 2017; Taqieddin et al., 2017; Soto et al., 2018; Angulo et al., 2020; Zhang and Lohse, 2023).
-Several experiments have been done, and they confirmed that bubble growth is mainly driven by mass transport due to the gradient of the dissolved hydrogen concentration.
-Glas and Westwater (1964) performed a simple experiment for hydrogen bubbles generated on a flat electrode.
-The experimental findings demonstrate that the asymptotic growth of electrochemically generated bubbles follows the same functional relationship given by the analytical solution for a suspended bubble growing in a supersaturated liquid, see Scriven (1959).
-Recent studies have already investigated this problem from different scales.
+Analytical solutions for this problem are not generally available, and only a few exact solutions can be derived using extremely simplified assumptions (Epstein1950; Scriven1959).
+However, they can still serve as valuable references for experimental and numerical studies (Glas1964; Brandon1985; dapkus1986; vanderlinde2017a; Taqieddin2017; Soto2018; Angulo2020; Zhang2023).
+Several experiments have confirmed that bubble growth is primarily governed by mass transport arising from gradients in the dissolved hydrogen concentration.
+Glas1964 performed a fundamental experiment for hydrogen bubbles generated on a flat electrode.
+The experimental findings demonstrate that the asymptotic growth of electrochemically generated bubbles follows the same functional relationship given by the analytical solution for a suspended bubble growing in a supersaturated liquid, see Scriven1959.
+Recent studies have already investigated this problem on different scales.
 Macroscopic processes are mainly convection phenomena in the electrolyte solution, which is observed at the electrolytic cell scale.
-Li et al.
-(2018) visualized experimentally on two-phase flow at the anode side of a proton exchange membrane electrolyzer.
-They understood that the inlet velocity does not affect bubble growth when the temperature and current density are constant.
-Moreover, they investigated gas bubble (hydrogen and oxygen) dynamics in a single-channel electrolyzer (Li et al., 2019).
-There, they showed that the bubble detachment diameter is inversely proportional to the flow velocity.
-Because of the strong influence of electrolyzer geometry on macro scale bubbly flow, Hreiz et al.
-(2015) studied several different electrochemical configuration sets.
-In the following work (Abdelouahed et al., 2014), Laser Doppler Velocimetry (LDV) measurements were conducted to observe bubble curtain behavior.
-Other experimental studies revealed that the bubble curtain velocity is a function of the average electric current (Hine and Murakami, 1980), the mass transfer coefficient is varied for horizontal and vertical electrodes (Fouad and Sedahmed, 1972), and bubble coverage is also a function of current density (Vogt and Balzer, 2005).
-Vogt (2012) has mentioned that when the bubble coverage tends to unity, the reaction is totally blocked.
-For the micro-scale, the microphysical phenomena surrounding individual bubbles constitute the overall behavior of the electrolysis process, and consequently, understanding the micro-scale bubble hydrodynamic is vital to reveal the governing mechanisms behind electrolysis.
-As the studied scale becomes smaller, however, the difficulty of the experiment rises dramatically.
-Several studies confirm that flow convection in the micro-area is caused by bubble growth, detachment, and coalescence (micro-convection) strongly influences the hydrogen diffusion boundary layer and changes mass transfer (Stephan and Vogt, 1979; Vogt and Balzer, 2005).
-van der Linde et al.
-(2018) used a specially designed electrode with pillars and pits to control micro-bubble formation, measuring the bubble radius and attachment force to verify the Fritz radius.
-Apart from bubble formation and growth, Bashkatov et al.
-(2024) investigated two modes of coalescence-induced bubble detachment using a dual platinum microelectrode system.
+li2018 visualized experimentally on two-phase flow at the anode side of a proton exchange membrane electrolyzer.
+They concluded that the inlet velocity does not affect bubble growth when the temperature and current density are constant.
+Furthermore, they examined hydrogen and oxygen bubble dynamics in a single-channel electrolyzer (li2019), revealing that the bubble detachment diameter varies inversely with flow velocity.
+Given the significant influence of electrolyzer geometry on macroscale bubbly flow, hreiz2015 analyzed a variety of electrochemical configurations.
+Abdelouahed2014 conducted Laser Doppler Velocimetry (LDV) measurements to observe the behavior of the bubble curtain.
+Other experimental studies revealed that the bubble curtain velocity is a function of the average electric current (hine1980), the mass transfer coefficient is varied for horizontal and vertical electrodes (fouad1972), and bubble coverage is also a function of current density (vogt2005).
+vogt2012 has mentioned that when the bubble coverage approaches unity, the reaction is totally blocked.
+For the micro-scale, the micro-physical phenomena surrounding individual bubbles govern the overall behavior of the electrolysis process, and consequently, understanding the micro-scale bubble hydrodynamics is vital to reveal the governing mechanisms behind electrolysis.
+As the studied scale becomes smaller, however, the difficulty of the experimentation increases dramatically.
+Several studies confirm that flow convection in the micro-area is caused by bubble growth, detachment, and coalescence.
+Such micro-convection strongly influences the hydrogen diffusion boundary layer and changes mass transfer (Stephan1979; vogt2005).
+For example, Linde2018 used a specially designed electrode with pillars and pits to control micro-bubble formation, measuring the bubble radius and attachment force to verify the Fritz radius.
+Apart from bubble formation and growth, Bashkatov2024 investigated two modes of coalescence-induced bubble detachment using a dual platinum microelectrode system.
 By combining high-speed imaging and electrochemical analysis, they demonstrated the importance of bubble-bubble interactions in the departure process.
 The impact of micro-convection induced by bubble dynamics on mass transfer is significant.
-Burdyny et al.
-(2017) demonstrated that micro-convection enhances mass transport primarily by reducing the departure diameter of bubbles from the electrode surface.
-With advanced visualization techniques, the gas concentration released in the liquid by bubbles can be investigated (Dani et al., 2007; Francois et al., 2011).
-However, experimental measurements are generally expensive and limited by the available measuring techniques, which usually provide global quantities (e.g., global mass transfer rate, rising velocities), and do not give information about local details, such as small scale multiphase flow field and local mass transfer rate.
+Burdyny2017 demonstrated that micro-convection enhances mass transport primarily by reducing the departure diameter of bubbles from the electrode surface.
+With advanced visualization techniques, the gas concentration released in the liquid by bubbles can be investigated (dani2007; francois2011).
+However, experimental measurements are generally expensive and constrained by the available measuring techniques, which usually provide global quantities (e.g., global mass transfer rate, rising velocities), and do not give information about local details, such as small-scale multiphase flow field and local mass transfer rate.
 Due to the presence of large amounts of gas, conventional optical techniques are unable to detect many critical aspects of multiphase flow fields.
-Computational Fluid Dynamics (CFD) has become an important tool for the comprehensive study of electrolytic complex multiphase flows (Hawkes et al., 2009; El-Askary et al., 2015).
+Computational Fluid Dynamics (CFD) has become an important tool for the comprehensive study of electrolytic complex multiphase flows (Hawkes2009; El-Askary2015).
 The significant increase in computational power has made direct numerical simulation an important alternative method for studying the detailed dynamics of mass transfer between fixed or deformable interfaces.
 The individual bubble growth, multiple bubbles coalescence, and detachment that occur in this region have been numerically studied for a long period.
-It is confirmed that the diffusion-driven bubble growth dominates the growth stage (Soto, 2019).
-The bubble detachment radius is approximated by the Fritz radius (Vogt et al., 2004).
-It can be derived by the force balance acting on the growing and detaching bubbles.
-Besides, the mass transfer behind the phenomenon still needs to be discussed.
-van der Linde et al.
-(2017); van der Linde et al.
-(2018) simulated the hydrogen concentration field around a growing hydrogen bubble in acidic electrolysis, using a body-fitted axisymmetric finite difference method.
-Sepahi et al.
-(2022) employed an Immersed Boundary Method (IBM) to simulate the mass transport for bubbles in gas-evolving electrolysis, indicating that the net transport within the system is governed by the effective buoyancy driving induced by the rising bubble.
+It is confirmed that the diffusion-driven bubble growth dominates the growth stage (Soto2018).
+The bubble detachment radius is approximated by the Fritz radius (vogt2004).
+It can be derived from the force balance acting on the growing and detaching bubbles.
+Besides, the mass transfer mechanisms underlying this phenomenon require further discussion.
+vanderlinde2017a; Linde2018 simulated the hydrogen concentration field around a growing hydrogen bubble in acidic electrolysis, using a body-fitted axisymmetric finite difference method.
+sepahi2022a employed an Immersed Boundary Method (IBM) to simulate the mass transport for bubbles in gas-evolving electrolysis, indicating that the net transport within the system is governed by the effective buoyancy driving induced by the rising bubble.
 Very few simulations of deformable-bubble growth, detachment, and rise in the context of electrolysis have been performed using modern sharp-interface methods such as VOF, level set, or front tracking.
-Moreover, some authors used VOF methods without mass transfer (Lafmejani et al., 2017).
-Most previous studies, such as those by Sepahi et al.
-(2022) and Khalighi et al.
-(2023), assumed undeformable spherical bubbles, neglecting bubble deformation that may arise due to lateral flow during detachment.
-Simulations of deformable bubbles without the spherical bubble assumption have not been performed very often due to the capillary time-step constraint.
-Thus, a rigid sphere approximation saves considerable CPU time.
-Although these models do not apply to the design of real systems, they still provide helpful information and reference solutions for the validation of other models.
-Recently, Gennari et al.
-(2022) developed a VOF-based phase-change model for diffusion-driven mass transfer problems using the one-fluid method and a novel algorithm to extrapolate the discontinuous velocity field across the interface to improve interface advection accuracy.
+Moreover, some authors used VOF methods without mass transfer (lafmejani2017).
+Most previous studies, such as those by sepahi2022a and khalighi2023, assumed undeformable spherical bubbles, neglecting bubble deformation that may arise due to lateral flow during detachment.
+Due to the time step limitations imposed by capillary effects, employing a rigid sphere approximation considerably reduces CPU time.
+Although these models are not suitable for designing actual systems, they still provide useful information and reference solutions for validating novel theories.
+Thus, simulations of bubble growth in electrolysis without the spherical bubble assumption have rarely been performed.
+Recently, Gennari2022 proposed a VOF-based phase-change model for diffusion-driven mass transfer problems using the one-fluid method and a novel algorithm to extrapolate the discontinuous velocity field across the interface to improve interface advection accuracy.
 This method is applied to study the growth of deformable bubbles on planar electrodes.
-In this paper, we investigate all these effects using direct numerical simulation (DNS), which means that we fully resolve all relevant scales of the hydrodynamic and concentration boundary layers.
-We take advantage of the free code repository Basilisk (http://basilisk.fr/).
+This study employs direct numerical simulation (DNS) to investigate all the aforementioned effects, meaning that we have fully resolved all relevant scales of the hydrodynamic and concentration boundary layers.
+Compared to existing numerical studies, we are the first to utilize three-dimensional DNS to simulate the deformable interface motion of hydrogen bubbles at the electrode.
+The fully resolved flow field provides crucial insight into the mechanisms of bubble growth, interaction, and detachment.
+All simulations are performed using the free code repository Basilisk (http://basilisk.fr/).
 2 Configuration and numerical methods
-To simplify the complicated mathematical description of a bubble growing at an electrode in the presence of a surrounding flow, consider a two-phase gas-liquid system represented in fig.
-1.
-We assume a constant room temperature, dilute liquid solutions ($\mathrm{KOH}$, 0.5 mol/L ), and no evaporation of water.
-Besides, in order to avoid further complications, self-ionisation of water is disregarded due to its low equilibrium constant at room temperature.
+To simplify the complicated mathematical description of a bubble growing at an electrode in the presence of a surrounding flow, consider a two-phase gas-liquid system represented in Fig.˜1.
+We assume a constant room temperature, dilute liquid solutions ($\mathrm{KOH}$, $0.5\mathrm{mol/L}$), and no evaporation of water.
+Besides, in order to avoid further complications, self-ionization of water is disregarded due to its low equilibrium constant at room temperature.
 The equation for the electric potential remains relatively simple.
@@
 2.1 Problem set-up
-The bubble growth and detachment are investigated first under axisymmetry.
-Due to the geometric axisymmetry, an axisymmetric simulation can also effectively address the problem when studying the growth of a single bubble on a circular electrode in addition to a full 3D simulation.
-A sketch of the axisymmetric setup and mesh grid is shown in fig.
-2.
-The axisymmetric condition is applied at the left boundary (Z axis), while an outflow boundary condition is set on the top boundary; the other boundaries are treated as no-slip wall.
+We start with axisymmetric simulations to validate the numerical mass transfer method provided by Gennari2022.
+Due to the geometric axisymmetric nature of the single bubble attached to a circular electrode, axisymmetric modeling not only effectively resolves the problem but also reduces computational costs while achieving a three-dimensional simulation.
+A sketch of the axisymmetric setup and mesh grid is shown in Fig.˜2.
+The symmetry axis is along $z$, while an outflow boundary condition is set on the top boundary.
+The other boundaries are treated as no-slip walls.
 The computational domain is a square.
-The domain’s characteristic length is $L_{0}=25L_{ref}$, where the reference length is the initial bubble diameter $D_{b}$.
-The bubble is initialized in simulations with the diameter $D_{b}=0.0127$ mm, The electrode is the flat end of a wire with a diameter of $D_{e}=10D_{b}=0.127$ mm, oriented along the axis of symmetry (Z axis).
-As for the more complex configuration, "mutiple nucleation sites", the geomertic axisymmetry does not exist anymore.
-To address the mutual interaction, a 3D configuration is needed.
-The multiple nucleation sites are equally spaced at the electrode center, as illustrated in fig.
-3.
-The initial liquid is set to be saturated, and the saturation ratio is $\zeta=c_{0}/c_{sat}=1$, where $c_{sat}$ is the hydrogen concentration in the saturated liquid and $c_{0}$ is the initial concentration of dissolved gas near the electrode surface.
-As mentioned in section 1, the production of dissolved gas at the electrode walls creates a locally supersaturated region ($\zeta>1$).
-It drives the growth of bubbles formed from microscopic pits on the electrode surface due to the heterogeneous nucleation, see Jones et al.
-(1999); van der Linde et al.
-(2017).
-According to the experiment, there should be no bubble present at initial time $t=0$.
-However, using the VOF method requires the volume fraction of gas need to be initialized.
+The domain’s size is $L_{0}=25D_{b}$, where $D_{b}$ is the initial bubble diameter.
+The bubble is initialized in simulations with the diameter $D_{b}=0.0127$ mm, The electrode is the flat end of a wire with a diameter of $D_{e}=10D_{b}=0.127$ mm, oriented along the radial axis (r axis).
+As for the more complex case of multiple nucleation sites, axial symmetry no longer holds.
+To address the mutual interaction between bubbles, a 3D configuration is required.
+The multiple nucleation sites are equally spaced at the electrode center, as illustrated in Fig.˜3.
+The initial liquid is set to be saturated, and the saturation ratio is $\zeta=c_{0}/c_{s}=1$, where $c_{s}$ is the hydrogen concentration in the saturated liquid and $c_{0}$ is the initial concentration of dissolved gas near the electrode surface.
+As mentioned in Section˜1, the production of dissolved gas at the electrode walls creates a locally supersaturated region ($\zeta>1$).
+It drives the growth of bubbles formed from microscopic pits on the electrode surface due to the heterogeneous nucleation, see Jones1999; vanderlinde2017a.
+According to the experiment, there should be no bubble present at the initial time $t=0$.
+However, using the VOF method requires the volume fraction of gas to be initialized.
 We wait for a nucleation time before computing the volume change.
@@
 The nucleation time varies due to many factors, like the electrode material and the current density.
-In present work, we set the nucleation time $t_{n}=0.02$ s, which is observed from Glas and Westwater (1964).
-The control parameters for the electrolytically generated bubbly flow are the cathodic current density, and the contact angles for different wettability of the electrode surface.
+In the present work, we set the nucleation time $t_{n}=0.02\,\mathrm{s}$, a value derived from experimental work(Glas1964).
+The control parameters for the electrolytically generated bubbly flow are the cathodic current density and the contact angles for different wettability of the electrode surface.
 The current density $I$ can give the molar flux of hydrogen ($\mathrm{H_{2}}$) by Faraday’s law,
@@
 where $I=i/A$, $i$ is the total electric current, $A$ is the cross-section area ($A=\pi/4D_{e}^{2}$); $F$ is Faraday’s constant ($F=96485.3$ As/mol).
-To account for the flux of H2 across the active area of the electrode, a Neumann boundary condition for the gas concentration is applied to the electrode wall (r-axis for axis-symmetric simulation, XY plane for 3D simulation),
+To account for the flux of H2 across the active area of the electrode, a Neumann boundary condition for the gas concentration is applied to the electrode wall (r-axis for axis-symmetric simulation, xy plane for three-dimensional simulation),
 $\frac{\partial c}{\partial z}$ $=\frac{J}{D}\qquad\text{for }r<\frac{D_{e}}{2},$ (3)
 $\frac{\partial c}{\partial z}$ $=0\qquad\text{for }r>\frac{D_{e}}{2}.$ (4)
-The boundary influx from the bubble-free region directly leads to an increase of the local concentration of dissolved hydrogen, leading to the development of a hydrogen concentration layer.
+The boundary influx from the bubble-free region directly leads to an increase in the local concentration of dissolved hydrogen, leading to the development of a hydrogen concentration layer.
 This process facilitates hydrogen bubble formation, driving its growth and detachment.
 Simulations performed in the present work are classified into two groups.
-The configurations are listed in Table 1.
-The axisymmetric configuration is mainly used for numerical method verification by comparing the bubble growth with analytical and experimental results.
-The 3D configuration is designed to study the bubble coalescence and detachment.
+The configurations are listed in Table˜1.
+The axisymmetric configuration is primarily employed for numerical verification and sensitivity analysis, whereas the 3D configuration is used for most of the simulations on bubble growth evolution.
 Figure 2: Sketch of the axisymmetric simulation setup.
@@
 Configuration
-Single bubble (Axisymmetric) $I(Am^{-2})$
-No. $\theta\degree$ $z\times r(L_{ref})$ mesh level 100
-1 90 $25\times 25$ 8 300
-2 90 $25\times 25$ 9 1000
-3 90 $25\times 25$ 10 -
-4 35 $25\times 25$ 8 -
-5 35 $25\times 25$ 9 -
-6 35 $25\times 25$ 10 -
-3D
-No. $\theta\degree$ $x\times y\times z(L_{ref})$ nucleation sites -
-7 90 $25\times 25\times 25$ 1 -
-8 90 $25\times 25\times 25$ 2 -
-9 90 $25\times 25\times 25$ 4 -
-10 35 $25\times 25\times 25$ 1 -
-11 35 $25\times 25\times 25$ 2 -
-12 35 $25\times 25\times 25$ 4 -
-Table 1: Simulation parameters for cases with a single bubble and mutiple bubble nucleation sites. All the simulation are performed at 3 different current densities, as listed in the last column.
-The physical parameters remain the same for all the simulation cases (see Table 2).
-The only exception is that gravity acceleration is set to zero for the axisymmetric configuration.
-To reduce the computational cost due to the time step limitation of the surface tension scheme, $\sigma$ is decreased by a factor of $10^{-4}$.
-Besides, the density of hydrogen is relatively small, and the large density ratio for this system ($\rho_{c}/\rho_{d}=12450$) would be a problem for the numerical simulation.
-It is observed that a large density ratio will slow down the convergence of the multigrid solver.
-To circumvent this problem, the hydrogen density is increased to $\rho_{d}=0.8\,\text{kg/m}^{3}$, the molar mass in increased by the same factor.
-so the volume change is not influenced ($\Delta V\propto M/\rho_{d}$).
-The zero gravity settings are only applied for the axisymmetric configuration.
-When it comes to the investigation of bubble detachment, gravity acceleration is applied at the direction of - Z in 3D configuration with the value of $9.8\mathrm{m/s^{2}}$.
+Single bubble (Axisymmetric)
+No. $\theta\degree$ $z\times r(D_{b})$ mesh level
+1 35–90 (sweep: 35, 40, …, 90) $25\times 25$ 8
+2 90(hydrophobic) $25\times 25$ 9
+3 90(hydrophobic) $25\times 25$ 10
+4 35(hydrophilic) $25\times 25$ 9
+5 35(hydrophilic) $25\times 25$ 10
+Three-dimensional
+No. $\theta\degree$ $x\times y\times z(D_{b})$ nucleation sites
+6 35–90 (sweep: 35, 40, …, 90) $25\times 25\times 25$ 1
+7 90(hydrophobic) $25\times 25\times 25$ 2
+8 90(hydrophobic) $25\times 25\times 25$ 4
+9 35(hydrophilic) $25\times 25\times 25$ 2
+10 35(hydrophilic) $25\times 25\times 25$ 4
+Table 1: Configurations for the three-dimensional and axisymmetric simulations.
+The physical parameters remain the same for all the simulation cases (see Table˜2). Gravity acceleration is applied in the z-z direction with the value of $9.8\mathrm{m/s^{2}}$.
 Symbol Properties Value Unit
-$\rho_{c}$ Electrolyte density 996 [$\mathrm{kg/m^{3}}$]
-$\rho_{d}$ Hydrogen density 0.8 [$\mathrm{kg/m^{3}}$]
-$g$ Gravity acceleration 9.8 [$\mathrm{m/s^{2}}$]
-$M$ Molar mass of hydrogen $2.0\times 10^{-2}$ [$\mathrm{kg/mol}$]
-$c_{0}$ H2 initial concentration 0.002 [$\mathrm{mol/m^{3}}$]
-$c_{\text{sat}}$ H2 saturated concentration 0.02 [$\mathrm{mol/m^{3}}$]
-$\mu_{c}$ Electrolyte viscosity $8.32\times 10^{-4}$ [$\mathrm{kg/(m\cdot s)}$]
-$\mu_{d}$ Hydrogen viscosity $8.96\times 10^{-6}$ [$\mathrm{kg/(m\cdot s)}$]
-$\nu_{c}$ Electrolyte kinematic viscosity $8.35\times 10^{-7}$ [$\mathrm{m^{2}/s}$]
-$\sigma$ Surface tension $7.5\times 10^{-2}$ [$\mathrm{N/m}$]
-$F$ Faraday’s constant $96485.3$ [$\mathrm{C/mol}$]
-$D$ Hydrogen diffusion coefficient $7.38\times 10^{-9}$ [$\mathrm{m^{2}/s}$]
-$L_{\text{ref}}$ Reference length $1.27\times 10^{-5}$ [$\mathrm{m}$]
-$D_{b}$ Initial bubble diameter $1.27\times 10^{-5}$ [$\mathrm{m}$]
-$D_{e}$ Electrode diameter $1.27\times 10^{-4}$ [$\mathrm{m}$]
+$\rho_{c}$ Electrolyte density 996 $\mathrm{kg/m^{3}}$]
+$\rho_{d}$ Hydrogen density 0.8 $\mathrm{kg/m^{3}}$]
+$g$ Gravity acceleration 9.8 $\mathrm{m/s^{2}}$]
+$M$ Molar mass of hydrogen 0.02 $\mathrm{kg/mol}$]
+$c_{0}$ H2 initial concentration 0.002 $\mathrm{mol/m^{3}}$]
+$c_{s}$ H2 saturated concentration 0.02 $\mathrm{mol/m^{3}}$]
+$\mu_{c}$ Electrolyte viscosity $8.32\times 10^{-4}$ $\mathrm{kg/(m\cdot s)}$]
+$\mu_{d}$ Hydrogen viscosity $8.96\times 10^{-6}$ $\mathrm{kg/(m\cdot s)}$]
+$\nu_{c}$ Electrolyte kinematic viscosity $8.35\times 10^{-7}$ $\mathrm{m^{2}/s}$]
+$\sigma$ Surface tension $7.5\times 10^{-6}$ $\mathrm{N/m}$]
+$F$ Faraday’s constant $96485.3$ $\mathrm{C/mol}$]
+$D$ Hydrogen diffusion coefficient $7.38\times 10^{-9}$ $\mathrm{m^{2}/s}$]
+$D_{b}$ Initial bubble diameter $1.27\times 10^{-5}$ $\mathrm{m}$]
+$R_{0}$ Initial bubble radius $6.35\times 10^{-6}$ $\mathrm{m}$]
+$D_{e}$ Electrode diameter $1.27\times 10^{-4}$ $\mathrm{m}$]
+$t_{n}$ Bubble nucleation time $0.02$ $\mathrm{s}$]
+$P_{0}$ Ambient pressure $1.01325\times 10^{5}$ $\mathrm{N/m^{2}}$]
+$T_{0}$ Ambient temperature $298.15$ $\mathrm{K}$]
+$\mathcal{R}$ Universal gas constant $8.3145$ $\mathrm{J\cdotp mol^{-1}\cdotp K^{-1}}$]
+$H_{s}^{cp}$ Henry’s law solubility constant for $\mathrm{H_{2}}$ $7.5\times 10^{-6}$ $\mathrm{mol/(m^{3}Pa)}$]
+$P$ Henry’s law partition coefficient 53.3 -
 Table 2: Physical properties in SI units.
@@
 The transport of hydrogen results in a surface-averaged concentration, $\bar{c_{e}}$ at the electrode surface.
-To illustrate the transport, we compare the $\bar{c_{e}}$ with the concentration value in the bulk liquid, for which we adopt the top of the domain.
-Since the electrolyte solution is initially saturated ($\zeta=1$), the initial value $c_{0}=c_{sat}$, it yields the Sherwood number of production of hydrogen,
-$\overline{\mathrm{Sh_{e}}}=\frac{JL_{ref}}{D(\bar{c_{e}}-c_{sat})}.$ (5)
+To illustrate the transport, we compare $\bar{c_{e}}$ with the bulk liquid concentration, defined here as the concentration at the top of the domain.
+Given that the electrolyte solution is initially saturated ($\zeta=1$), and $c_{0}=c_{s}$, this yields the following Sherwood number for hydrogen production,
+$\overline{\mathrm{Sh_{e}}}=\frac{JD_{b}}{D(\bar{c_{e}}-c_{s})}.$ (5)
 The bar symbol is used to mark the surface-averaged response parameters.
-By introducing the boundary layer thickness $\delta_{\mathrm{H_{2}}}=D\Delta c/J$, this Sherwood number could be expressed as $\mathrm{Sh}=D_{b}/\delta_{\mathrm{H_{2}}}$.
-For pure diffusion, the $\delta_{\mathrm{H_{2}}}$ ultimately reaches the electrolysis cell height irrespective of the current density so that the Sherwood could maintain stability for all the cases without the effect of the bubbles.
-For the mass transfer considering local volume change, the hydrogen transport to the bubble will be considered. This leads to the expression for the bubble Sherwood number,
-[…]
-where $A_{\Sigma}$ is the interface surface area, $4\pi R$, and the concentration difference between the electrode and bubble interface, $c_{e}-c_{\Sigma}$.
-According to eq.
-31, the concentration of the interface $c_{\Sigma}=c_{sat}$.
-Since we assume the configuration with an ambient temperature and pressure, the mass transfer rate across the interface can be defined as:
-[…]
-where $\mathcal{R},T_{0},P_{0}$ are the universal gas constant, ambient temperature, and pressure, respectively. Then it leads to a rather simple expression of bubble Sherwood number,
-$\overline{\mathrm{Sh_{e}}}=\frac{P_{0}}{\mathcal{R}T_{0}}\frac{2R}{D(\bar{c_{e}}-c_{\Sigma})}\frac{dR}{dt}$ (8)
-Apart from this crucial Sherwood number, which measures the mass transfer, other nondimensional numbers are also vital as control parameters for physics.
-First, the Schmidt number compares momentum diffusion to mass diffusion,
+Introducing the boundary layer thickness $\delta_{\mathrm{H_{2}}}=D(\bar{c_{e}}-c_{s})/J$, this Sherwood number can be expressed as $\overline{\mathrm{Sh_{e}}}=D_{b}/\delta_{\mathrm{H_{2}}}$.
+The boundary layer thickness is, in principle, time dependent, so the Sherwood number also is time dependent.
+However, after several bubble detachments or in the presence of externally forced flow, the boundary layer thickness would eventually reach a statistical steady state, and a statistically steady Sherwood number will be reached.
+For mass transfer accounting for local volume change, the Sherwood number of the bubble is expressed as follows to quantify the hydrogen transport across the interface,
+$\overline{\mathrm{Sh_{b}}}=\frac{D_{b}\int_{\Sigma}\dot{m}\,ds}{A_{\Sigma}MD(\bar{c_{e}}-c_{s})},$ (6)
+where $A_{\Sigma}=4\pi R^{2}$ is the interface surface area for a spherical bubble.
+And $\bar{c_{e}}-c_{s}$ represents the concentration difference between the electrode and bubble interface as the concentration of the interface $c_{\Sigma}=c_{s}$.
+Since we assume the configuration with an ambient temperature and pressure, the mass transfer rate across the interface can be defined as
+$\int_{\Sigma}\dot{m}\,ds=\left(\frac{P_{0}}{\mathcal{R}T_{0}}\right)4\pi R^{2}\frac{dR}{dt}M,$ (7)
+where $\mathcal{R},T_{0},P_{0}$ are the universal gas constant, ambient temperature, and pressure, respectively. Then it leads to a rather simple expression of the bubble Sherwood number,
+$\overline{\mathrm{Sh_{b}}}=\frac{P_{0}}{\mathcal{R}T_{0}}\frac{2R}{D(\bar{c_{e}}-c_{\Sigma})}\frac{dR}{dt}.$ (8)
+Apart from the Sherwood number, which measures the mass transfer, other important nondimensional numbers are the Schmidt, Galileo, and Bond numbers.
+First, the Schmidt number compares momentum diffusion to mass diffusion and is defined as
 $\mathrm{Sc}=\frac{\nu_{c}}{D},$ (9)
 where the $\nu_{c}$ is the kinematic viscosity of the electrolyte.
-For these configurations with gravity added, the dynamics of bubbles in a gravitational field can be described by the Galilei and Bond numbers.
-The Galilei number compares the gravitational and viscous forces and is defined as:
+The ratios of dominant forces acting on the bubble can be described using two additional dimensionless numbers: the Galileo and Bond numbers.
+The Galileo number, which compares gravitational to viscous forces, is given by
 $\mathrm{Ga}=\sqrt{\frac{\rho_{c}gD_{b}^{2}}{\nu_{c}^{2}}}.$ (10)
-The ratio of gravitational and capillary forces is estimated by the Bond number,
+When the bubble approaches detachment, gravitational effects become dominant and can be quantified by the Bond number, which represents the ratio of gravitational to capillary forces
 $\mathrm{Bo}=\frac{\rho_{c}gD_{b}^{2}}{\sigma}.$ (11)
-2.3 Governing equation
-A model of the microscopic process involves the solution of the two-phase incompressible Navier-Stokes equations with phase change, surface tension, gravity, and contact line dynamics on the wall.
+2.3 Fluid dynamical equations
+A model of the microscopic process involves solving the two-phase incompressible Navier-Stokes equations with phase change, surface tension, gravity, and contact line dynamics on the wall.
 Due to the constant temperature assumption, no equation for the thermal energy is needed.
-The gas phase is called $\Omega_{d}(t)$, for the disperse phase, and the liquid phase is called $\Omega_{c}(t)$, for the continous phase.
+The gas phase is called $\Omega_{d}(t)$, for the disperse phase, and the liquid phase is called $\Omega_{c}(t)$, for the continuous phase.
 These two subdomains are separated by an infinitely thin interface $\Sigma(t)$.
@@
 $\nabla\cdot\mathbf{u}$ $=0\quad\text{in }\Omega\backslash\Sigma,$ (12)
-$\begin{split}\partial_{t}(\rho\mathbf{u})+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u})&=-\nabla p+\nabla\cdot(2\mu\mathbf{D})\\ &\quad+\rho\mathbf{g}\quad\text{in }\Omega\backslash\Sigma.\end{split}$ (13)
+$\begin{split}\partial_{t}(\rho\mathbf{u})+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u})&=-\nabla p+\nabla\cdot(2\mu\mathbf{D})\\ &\quad+\rho\mathbf{g}\quad\text{in }\Omega\backslash\Sigma,\end{split}$ (13)
 where the density $\rho$ and viscosity $\mu$ remain constant in $\Omega_{d}$ and $\Omega_{c}$.
-The eq.
-12 is the continuity equation, where $\mathbf{u}$ represents the velocity field.
-In the balance of momentum eq.
-13, $p$ is the static pressure, $\mathbf{D}$ is the deformation tensor, and $g$ represents body force, which is gravity acceleration in this system.
-So, $a$ is replaced by $g$ for the following sections.
-The eq.
-12 and eq.
-13 are valid everywhere in the domain except at the interface, where additional conditions are needed (Tryggvason et al., 2011).
+The Eq.˜12 is the continuity equation, where $\mathbf{u}$ represents the velocity field.
+In the balance of momentum Eq.˜13, $p$ is the static pressure, $\mathbf{D}$ is the deformation tensor, and $g$ represents body force, which is the gravitational acceleration in this system.
+So, $a$ is replaced by $g$ in the following sections.
+The Eq.˜12 and Eq.˜13 are valid everywhere in the domain except at the interface, where additional conditions are needed (Tryggvason_Scardovelli_Zaleski_2011).
 The continuity equation requires that the amount of mass that leaves one phase $\Omega_{d}(\Omega_{c})$ must be transferred to another phase $\Omega_{c}(\Omega_{d})$ since the infinitely thin interface region can not store any mass.
 It results in a jump condition across the interface,
-$\rho(\mathbf{u}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}=0,$ (14)
-where the jump notation has been introduced; $\mathbf{u_{\Sigma}}$ is the interface velocity and $\dot{m}$ is the mass transfer rate $\mathrm{kg/m^{2}s}$.
-The second jump condition is derived by applying the conservation of momentum to a control volume with an infinitely small thickness around the interface, and it reads,
-$\mathbf{u}\otimes(\mathbf{u}-\mathbf{u_{\Sigma}})+p\mathbf{I}-2\mu\mathbf{D}\cdot\mathbf{n_{\Sigma}}=\sigma k\mathbf{n_{\Sigma}}+\nabla_{\Sigma}\sigma.$ (15)
+$\|\rho(\mathbf{u}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}\|=\|\dot{m}\|=0,$ (14)
+where the jump notation has been introduced(e.g.
+$\|\rho\|=\rho_{c}-\rho_{d}$); $\mathbf{u_{\Sigma}}$ is the interface velocity and $\dot{m}$ is the mass transfer rate $\mathrm{kg/m^{2}s}$.
+The second jump condition is derived by applying the conservation of momentum to a control volume with an infinitely small thickness around the interface, and it reads
+$\|\rho\mathbf{u}\otimes(\mathbf{u}-\mathbf{u_{\Sigma}})+p\mathbf{I}-2\mu\mathbf{D}\|\cdot\mathbf{n_{\Sigma}}=\sigma k\mathbf{n_{\Sigma}}+\nabla_{\Sigma}\sigma,$ (15)
 where $\mathbf{I}$ is the unit tensor, $\sigma$ is the surface tension and $k$ is the curvature of the interface.
-In the problem of hydrogen bubble growth, the interface is considered to have uniform surface tension, besides a no-slip boundary condition is applied at the interface, eq.
-15 can be further simplified, see Fleckenstein and Bothe (2015).
-$\left(p\mathbf{I}-2\mu\mathbf{D}\|\cdot\mathbf{n_{\Sigma}}\right)=\sigma k\mathbf{n_{\Sigma}}$ (16)
-The numerical method used in this work for interface transport is the Volume of Fluid (VOF) method, and is combined with one-fluid formulation of the governing equations.
-In the one-fluid approach, the jump condition eq.
-14 and eq.
-16 are replaced by source terms that act at the interface as singularities ($\delta$ function), a single set of Navier-Stokes equations for the entire domain $\Omega$ is solved.
+In the problem of hydrogen bubble growth, the interface is considered to have uniform surface tension; besides, a no-slip boundary condition is applied at the interface.
+Eq.˜15 can be further simplified fleckenstein2015 in the form
+$\|p\mathbf{I}-2\mu\mathbf{D}\|\cdot\mathbf{n_{\Sigma}}=\sigma k\mathbf{n_{\Sigma}}.$ (16)
+The numerical method used in this work for interface transport is the Volume of Fluid (VOF) method, and is combined with a one-fluid formulation of the governing equations.
+In the one-fluid approach, the jump conditions Eq.˜14 and Eq.˜16 are replaced by source terms that act at the interface as singularities ($\delta$ function), a single set of Navier-Stokes equations for the entire domain $\Omega$ is solved as follows
 $\nabla\cdot\mathbf{u}$ $=\dot{m}\left(\frac{1}{\rho_{d}}-\frac{1}{\rho_{c}}\right)\delta_{\Sigma},$ (17)
 $\partial_{t}\mathbf{u}+\nabla\cdot(\mathbf{u}\otimes\mathbf{u})$ $=\frac{1}{\rho}\left[-\nabla p+\nabla\cdot(2\mu\mathbf{D})\right]+\frac{\sigma k\mathbf{n_{\Sigma}}}{\rho}\delta_{\Sigma},$ (18)
-where the $\delta_{\Sigma}$ is the surface Dirac function which has a nonzero value only at the interface.
-So the system of eq.
-17 eq.
-18 is valid in the whole domain.
+where the $\delta_{\Sigma}$ is the surface Dirac distribution.
+Therefore, the system of Eqs˜17 and 18 is valid for the whole domain.
 Then, to determine the location of the interface, a marker function is required.
-The Heaviside function serves this purpose,
+The Heaviside function serves this purpose
 $H(x,t)=\begin{cases}1,&\text{if }x\in\Omega_{c},\\ 0,&\text{if }x\in\Omega_{d}.\end{cases}$ (19)
-The transport equation for Heaviside function $H(x,t)$ can be obtained from the following integral balance for a control volume $V$.
-[…]
+The transport equation for Heaviside function $H(x,t)$ can be obtained from the following integral balance for a control volume $V$
+$\begin{split}\int_{V}\partial_{t}H\,dV+\oint_{\partial V}H\mathbf{u}\cdot\mathbf{n}\,dS\\ +\int_{\Sigma}(\mathbf{u}_{c}-\mathbf{u}_{\Sigma})\cdot\mathbf{n}_{\Sigma}\,dS&=0,\end{split}$ (20)
 where the second term on the LHS represents the convective transport, and the last term is a source term that accounts for the mass transfer across the interface, which is null when $\dot{m}=0$.
-Converting the surface integral to volume integral, we can write it into differential form,
+Converting the surface integral to a volume integral, we can write it in differential form,
 $\partial_{t}H+\nabla\cdot(H\mathbf{u})+\frac{\dot{m}}{\rho_{c}}\delta_{\Sigma}=0.$ (21)
-eq. 21 is then served to compute the volume fraction, see section 2.5.
+Eq.˜21 is then used to compute the volume fraction, see Section˜2.5.
 2.4 Concentration transport equation
-This model can be used to compute the concentration field of the soluble hydrogen in a two-phase system with mass transfer by applying the two-scalar method of Fleckenstein and Bothe (2015).
-In the present study, we focus on pure incompressible gas bubbles, and we assume that no electrolyte species exists in the disperse phase (i.e., the electrolyte is not volatile).
-The system contains two different species: hydrogen (denoted with subscript 1) and electrolyte liquid (denoted with subscript 2).
-The overall mass transfer is entirely given by the transfer of the hydrogen species.
+This model can be used to compute the concentration field of the soluble hydrogen in a two-phase system with mass transfer by applying the two-scalar method of fleckenstein2015.
+In the present study, we focus on pure incompressible gas bubbles, and we assume that no electrolyte species exists in the gas phase (Phase (i.e., the electrolyte is not volatile).
+The system contains two species: hydrogen (denoted by subscript 1) and electrolyte liquid (denoted by subscript 2).
+Overall mass transfer is entirely governed by the transport of the hydrogen species.
 It is worth pointing out that we do not need to solve the mass balance in the dispersed phase since no mixture exists inside the bubbles (pure hydrogen gas).
-The mass balance of hydrogen in the continous side $\Omega_{c}$ reads
-$\partial_{t}\rho^{1}+\nabla\cdot(\rho^{1}\mathbf{u^{1}})=R,$ (22)
-where $\rho^{1}$ is the partial density of hydrogen, $\mathbf{u^{1}}$ denotes the hydrogen velocity.
-$R$ is the reaction term that can model chemical reactions.
-It is coupled with the jump condition for the conservation of mass,
-$\|\rho^{1}(\mathbf{u^{1}}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}||=||\dot{m}^{1}||=0.$ (23)
-The average phase density and velocity of are derived from the respective species terms,
+The mass balance of hydrogen in the continuous phase domain $\Omega_{c}$ reads
+$\partial_{t}\rho^{1}+\nabla\cdot(\rho^{1}\mathbf{u^{1}})=0,$ (22)
+where $\rho^{1}$ is the partial density of hydrogen and $\mathbf{u^{1}}$ denotes the hydrogen velocity. It is coupled with the jump condition for the conservation of mass
+$\|\rho^{1}(\mathbf{u^{1}}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}\|=\|\dot{m}^{1}\|=0.$ (23)
+The average phase density and velocity are derived from the respective species terms
 $\rho_{c}=\rho^{1}+\rho^{2},$ (24)
@@
 $\rho_{c}\mathbf{u}=\rho^{1}\mathbf{u^{1}}+\rho^{2}\mathbf{u^{2}},$ (25)
-where the superscript $2$ denotes the species electrolyte. Then the transport equation of hydrogen for incompressible flow can be written as
-$\partial_{t}\rho^{1}+\mathbf{u}\cdot\nabla\rho^{1}+\nabla\cdot J^{1}=R,$ (26)
+where the superscript $2$ denotes the electrolyte species. Then the transport equation of hydrogen for incompressible flow can be written as
+$\partial_{t}\rho^{1}+\mathbf{u}\cdot\nabla\rho^{1}+\nabla\cdot J^{1}=0,$ (26)
 and the diffusive flux of hydrogen is
 $J^{1}=\rho^{1}(\mathbf{u^{1}}-\mathbf{u}).$ (27)
-The mass transfer rate of hydrogen can be derived from eq. 23,eq. 14.
-$\begin{split}\dot{m}^{1}&=\rho^{1}(\mathbf{u}-\mathbf{u}_{\Sigma})\cdot\mathbf{n}_{\Sigma}+\rho^{1}(\mathbf{u}^{1}-\mathbf{u})\cdot\mathbf{n}_{\Sigma}\\ &=\frac{\rho^{1}}{\rho_{c}}\cdot\dot{m}+J^{1}\cdot\mathbf{n}_{\Sigma}\end{split}$ (28)
+The mass-transfer rate of hydrogen can be derived from Eqs˜23 and 14
+$\begin{split}\dot{m}^{1}&=\rho^{1}(\mathbf{u}-\mathbf{u}_{\Sigma})\cdot\mathbf{n}_{\Sigma}+\rho^{1}(\mathbf{u}^{1}-\mathbf{u})\cdot\mathbf{n}_{\Sigma}\\ &=\frac{\rho^{1}}{\rho_{c}}\cdot\dot{m}+J^{1}\cdot\mathbf{n}_{\Sigma},\end{split}$ (28)
 which shows that the mass transfer contains both a convective term and a diffusive term.
@@
 $J^{1}=-D^{1}\nabla\rho^{1},$ (29)
-where the $D^{1}$ is the hydrogen diffusion coefficient. Combining the eq. 29 and eq. 28, the mass transfer rate of hydrogen reads
-$\dot{m}=-\frac{M^{1}D^{1}}{1-\frac{\rho^{1}}{\rho_{c}}}\frac{\partial c^{1}}{\partial\mathbf{n_{\Sigma}}},$ (30)
+where the $D^{1}$ is the hydrogen diffusion coefficient. Combining Eqs (29) and (28), the mass-transfer rate of hydrogen reads
+$\dot{m}=-\frac{M^{1}D^{1}}{1-\frac{\rho^{1}}{\rho_{c}}}\frac{\partial c^{1}}{\partial n_{\Sigma}},$ (30)
 where the molar concentration has been introduced, i.e.
-$c^{1}=\rho^{1}/M^{1}$ and $M^{1}$ is the molar mass.
-It gives the mass transfer rate evaluated from the continuous side of the interface, where the concentration of the dissolved species is generally variable.
-Since we have only hydrogen, which can be transferred between phases, in the following, we will omit the species indicator as we will always refer to the concentration of the soluble hydrogen in the liquid phase $\Omega_{c}$.
+$c^{1}=\rho^{1}/M^{1}$ with $M^{1}$ the molar mass.
+It provides the mass-transfer rate evaluated from the continuous side of the interface, which corresponds to $n_{\Sigma}=1$.
+Since only hydrogen can be transferred between phases, we will omit the species indicator in what follows.
+We will only refer to the concentration of soluble hydrogen in the liquid phase domain $\Omega_{c}$.
 To compute the concentration gradient, we need the hydrogen concentration at the liquid side of the interface.
-For a gas-liquid system at equilibrium, we can employ Henry’s law to compute the concentration on the liquid side of the interface.
-$(c_{c})_{\Sigma}=\frac{(c_{d})_{\Sigma}}{\mathrm{He}},$ (31)
-where $\mathrm{He}$ is the Henry coefficient, and it is taken as a constant for the present work.
+For a gas-liquid system at equilibrium, we can employ Henry’s law to compute the concentration on the liquid side of the interface
+$(c_{c})_{\Sigma}=\frac{(c_{d})_{\Sigma}}{P},$ (31)
+where $P$ is the partition coefficient, which can be deduced from $P=1/({\mathcal{R}T_{0}H_{s}^{cp}})$, and is taken as a constant for the present work.
 So, the hydrogen concentration at the liquid side of the interface $(c_{c})_{\Sigma}$ is immediately computed.
-$(c_{d})_{\Sigma}=\rho_{d}/M$ is a constant as the density should be equal everywhere inside the bubble ($\Omega_{d}$).
-For numerical integration, we rewrite the mass balance eq.
-26 for the hydrogen diffusion process.
+$(c_{d})_{\Sigma}=\rho_{d}/M$ is a constant, as the density should be constant everywhere inside the bubble ($\Omega_{d}$).
+For numerical integration, we rewrite the mass balance Eq.˜26 for the hydrogen diffusion process.
 As we discussed, the mass balance will be done in the continuous region $\Omega_{c}$ and for a non-reactive flow ($R=0$).
-Since the hydrogen dissolved in $\Omega_{c}$ is the only species that we need for mass transportation.
-The phase indicator will be omitted in the following, i.e., $c=c_{c}$.
-[…]
-Combine the jump condition eq. 23, the final differential form could be deduced.
+Since the hydrogen dissolved in $\Omega_{c}$ is the only species that we need for mass transportation
+$\begin{split}\int_{V}\partial_{t}c\,dV&+\oint_{\partial V}(c\mathbf{u}-D\nabla c)\cdot\mathbf{n}\,dS\\ &+\int_{\Sigma}c(\mathbf{u}-\mathbf{u}_{\Sigma})\cdot\mathbf{n}_{\Sigma}\,dS\\ &=0,\end{split}$ (32)
+The phase indicator will be omitted in the following, i.e., $c=c_{c}$. Combine the jump condition Eq.˜23, and the final differential form could be deduced,
 $\partial_{t}c+\mathbf{u}\cdot\nabla c=\nabla\cdot\left(D\nabla c\right)-\frac{\dot{m}}{M}\delta_{\Sigma}.$ (33)
 2.5 Numerical methodology
-The phase change model is derived from the work of Gennari et al.
-(2022).
-The governing equation shown in section 2.3 is solved by open source solver Basilisk (HTTP://basilisk.fr/) which is a finite-volume P.D.E.
-solver on adaptive Cartesian grids.
-Using adaptive mesh refinement (AMR) in regions with large gradients makes the approach particulally suitable for multiscale processes such as interfacial flows.
+The phase change model is derived from the work of Gennari2022.
+The governing equation shown in Section˜2.3 is solved using the free open science platform Basilisk (HTTP://basilisk.fr/), which provides finite-volume partial differential equation (P.D.E.) solvers on adaptive cartesian grids.
+Using quadtree/octree adaptive mesh refinement (AMR) in regions with large gradients makes the approach particularly suitable for multiscale processes such as interfacial flows.
 In interfacial flows, a fine mesh is typically required around the gas-liquid interface but not the entire domain.
-The shape of the domain is always a square $L_{0}\times L_{0}$ in axisymmetric configuration (cube $L_{0}\times L_{0}\times L_{0}$ in 3D).
-The grid is organized following a hierarchical quadtree (octree) structure, where each cell can be further divided into four child cells (eight in 3D), and a level is assigned to each cell according to the relative tree structure.
-The root cell is at level 0, and its size $\Delta$ is the same as the whole numerical domain ($\Delta=L_{0}$).
-A generic cell at level $l$ has a resolution of $\Delta(l)=\frac{L_{0}}{2^{l}}$.
+The shape of the domain is always a square $L_{0}\times L_{0}$ in the axisymmetric configuration (a cube $L_{0}\times L_{0}\times L_{0}$ in three dimensions).
+The grid is organized following a hierarchical quadtree/octree structure, where each cell can be further divided into four child cells (eight in three dimensions), and a level is assigned to each cell according to its position in the tree structure.
+The root cell is at level 0, and its size $\Delta$ is the same as that of the whole numerical domain ($\Delta=L_{0}$).
+A generic cell at level $l$ has size $\Delta(l)=\frac{L_{0}}{2^{l}}$.
 The grid structure in Basilisk allows neighboring cells to vary by up to one level, meaning each cell edge/face can communicate with no more than two finer edges/faces.
-An example of a quadtree grid with different levels of refinement and the relative tree structure is shown in fig.
-4.
-Figure 4: An example of quadtree discretisation (left) together with its logical representation (right). The level of the cells in the tree is also given (Popinet, 2012).
 The VOF method is one of the most widely used numerical approaches for the modeling of two-phase immiscible fluids.
-The starting point for the derivation of the VOF approach is the one-fluid formulation presented in section 2.3, which is vital to the phase change problem.
+The starting point for the derivation of the VOF approach is the one-fluid formulation presented in Section˜2.3, which is vital to the phase change problem.
 The volume fraction of the continuous phase is defined as
+$f_{c}$ $=\frac{1}{V}\int_{V}H\,dV,$ (34)
+and the value of $f_{c}$ is within the set $[0,1]$, depending on the amount of liquid in the cell.
 […]
-and the value of $f_{c}$ is within the set $[0,1]$, depending on the amount of liquid in the cell.
-$f_{c}=\begin{cases}0&\text{if the cell is pure gas},\\ 1&\text{if the cell is pure liquid},\\ [0,1]&\text{if the cell is mixed}.\end{cases}$ (35)
 The volume fraction of the dispersed phase $f_{d}$ is implicitly described by the relationship $f_{c}+f_{d}=1$.
 So, only one transport equation of H needs to be solved.
-By applying the incompressibility constraint, the integrated form of eq.
-21 is
-[…]
-According to the finding of Scardovelli and Zaleski (1999), the integration of the transport eq.
-36 is performed in two steps, namely the reconstruction step and the propagation step.
+By applying the incompressibility constraint, the integrated form of Eq.˜21 is
+$\begin{split}\frac{\partial}{\partial t}\int_{V}H\,dV&+\frac{1}{V}\int_{V}\nabla\cdot(H\mathbf{u})\,dV\\ &+\frac{1}{V}\int_{V}\frac{\dot{m}}{\rho_{c}}\delta_{\Sigma}\,dV=0.\end{split}$ (36)
+The integration of the transport Eq.˜36 is performed in two steps, namely the reconstruction step and the propagation step.
 First, the interface is approximated with a line/plane in each interfacial cell.
-Second, the fluxes of volume fraction across the cell boundaries are computed, and eq.
-36 is integrated in time.
-The geometric reconstruction of the interface is based on the piecewise linear interface construction (PLIC) method, where the interface is approximated as a line (plane) in axisymmetric configuration (3D).
+Second, the fluxes of volume fraction across the cell boundaries are computed, and Eq.˜36 is integrated in time.
+The geometric reconstruction of the interface is based on the piecewise linear interface construction (PLIC) method, where the interface is approximated as a line (plane) in an axisymmetric (three-dimensional) configuration.
 3 Implementation
-In this section, we show the simulation results of several different configurations; as described in section 2.1.
-This numerical study is designed to simulate the experiments conducted by Glas and Westwater (1964).
-The experimental result confirm the driven force of bubble growth is diffusion and measure the bubble growth rate.
-The comparsion between following numerical and experimental growth rate will verify the accuracy of simulation.
-The physical properties corresponding to an alkaline solution (typically used in industrial water electrolysis) are shown in Table 2.
-The simulation results are presented with dimensional units for comparison with experimental measurements.
+In this section, we show the simulation results for several different configurations, as described in Section˜2.1.
+This numerical study is designed to simulate the experiments conducted by Glas1964.
+The experimental results confirm that the driving force for bubble growth is diffusion and measure the bubble growth rate.
+The physical properties corresponding to an alkaline solution (typically used in industrial water electrolysis) are shown in Table˜2.
+It should be noted that the original surface tension $\sigma=0.075\mathrm{N/m}$ is decreased by a factor of $10^{-4}$ and the hydrogen density $\rho_{d}=0.08\,\text{kg/m}^{3}$ is increased to $\rho_{d}=0.8\,\text{kg/m}^{3}$.
+The molar mass is scaled by the same factor, ensuring that the relative volume change remains unaffected ($\Delta V\propto M/\rho_{d}$).
+Sensitivity tests are therefore required to assess the effects of the decreased surface tension and increased hydrogen density.
+Decreasing the surface tension is introduced to reduce computational cost, as the surface-tension scheme imposes a restrictive time-step constraint.
+A comparison of simulations with modified $\sigma$ in Fig.˜4 indicates that increasing surface tension has only a minor influence on bubble growth.
+An analysis of the Weber number for different values of $\sigma$ further shows that surface tension remains the dominant force during the bubble growth stage, even after $\sigma$ is decreased.
+Bubble detachment, however, is significantly more sensitive to the choice of $\sigma$, as reflected by the Bond number.
+To maintain consistency with the theoretical prediction of the detachment radius, the same decreased value of $\sigma$ is adopted.
+Regarding hydrogen density, its physical value is relatively small, leading to a large density ratio $(\rho_{c}/\rho_{d}=12450)$, which would be a problem for numerical simulations.
+A high density ratio is observed to slow the convergence of the multigrid solver.
+Sensitivity tests for different hydrogen densities, shown in Fig.˜4, demonstrate that increasing $\rho_{d}$ has a negligible effect on bubble growth.
+Based on the results above, we can confirm that a decrease of surface tension $\sigma$ by a factor of $10^{-4}$ and an increase of hydrogen density to $10\times\rho_{d}$ provide an acceptable balance between numerical stability, computational efficiency and physical accuracy.
+Figure 4: Sensitivity test on decreased surface tension (a), and increased hydrogen density (b).
+The simulation results in this work are presented in dimensional units to facilitate comparison with experimental measurements.
 According to the analytical and experimental work, the growth of electrochemically generated bubbles follows the same functional relationship as the solution for bubble growth in a supersaturated liquid described by the Scriven model.
-Different stages or regimes for bubble growth have been characterized by a power law: growth controlled by inertia ($R\propto t$), by diffusion ($R\propto t^{1/2}$) or by reaction limitation ($R\propto t^{1/3}$) , see (Angulo et al., 2020).
+Different stages or regimes for bubble growth have been characterized by a power law: growth controlled by inertia ($R\propto t$), by diffusion ($R\propto t^{1/2}$) or by reaction limitation ($R\propto t^{1/3}$) (Angulo2020).
 The growth behavior ($R\propto t^{1/2}$) reflects a standard analytical solution for diffusive bubble growth.
-To validate the specific relationship, we perform a series of axisymmetric simulations first.
+To validate the specific relationship, we first perform a series of axisymmetric simulations on bubble growing at the contact angle of $\theta=90\degree$.
 3.1 Axisymmetric simulation validation
-In this section, we assume there is only one bubble (with an initial diameter $D_{b}$) on the cathode.
-In fig.
-5, the left-hand side of the snapshots illustrates hydrogen evolution during the bubble’s growth stage, where color represents the dissolved hydrogen concentration in the liquid phase.
-For visualization consistency, the gas phase inside the bubble will also be colored, but will not be calculated (based on the assumption of constant pressure inside the bubble) and will remain at 1.
-The same map is applied to the rest of the snapshots in the following.
-Hydrogen is produced at the electrode surface, so hydrogen saturation near the electrode wall is more significant than that in the bulk liquid.
-Completing the visualization, the right side of the snapshots displays the symmetric grid cell and interface, clearly demonstrating the mesh refinement near the gas-liquid interface.
-The growth rate of a single bubble confirms the result of the diffusion-controlled power law ($R\propto t^{1/2}$).
-As shown in fig.
-6, a mesh independence study demonstrates that mesh refinement level 8 will be enough for interface tracking.
-A logarithmic plot of the bubble growth for different current densities between $t=0.02\mathrm{s}$ and $t=0.2\mathrm{s}$ of our simulation (mesh refinement level 8) is shown in fig.
-7.
+In this Section, we assume there is a single bubble (with an initial diameter $D_{b}$) on the cathode.
+In Fig.˜5, the left-hand side of the snapshots illustrates hydrogen evolution during the bubble’s growth stage, where color represents the dissolved hydrogen concentration in the liquid phase.
+For visualization consistency, the gas phase inside the bubble will also be colored, but will not be calculated (based on the assumption of constant pressure inside the bubble) and will remain at $1\mathrm{mol/m^{3}}$.
+The same color map is applied to the rest of the snapshots in the following.
+Hydrogen is produced from the electrode surface, and the saturation near the electrode wall is more significant than that in the bulk liquid.
+Completing the visualization, the right side of the snapshots displays the grid and the moving interface, clearly showing adaptive quadtree mesh refinement.
+A mesh independence study has been performed to demonstrate that a mesh refinement level of 8 will be sufficient for accurate resolution.
+The following logarithmic plot Fig.˜6 presents bubble growth for five different current densities between $t_{n}=0.02\,\mathrm{s}$ and $t=0.2\,\mathrm{s}$.
 In our simulation case, the hydrogen flux is applied before the bubble nucleation time ($t_{n}=20\mathrm{ms}$), with the average hydrogen concentration on the electrode wall increasing in time.
 After nucleation, the concentration difference drives bubble growth through the diffusion process.
-As a result, the bubble growth behavior shows some discrepancies from the analytical solution during the early stages of diffusion.
-For smaller current density $I=100\mathrm{A/m^{2}}$, the growth exponent is smaller than $1/2$.
-For much larger current density $I=1000\mathrm{A/m^{2}}$, the growth follows the power law ($R\propto t^{0.8}$) better.
-In a similar simulation using the immersed boundary method, see Khalighi et al.
-(2023), the growth exponent 0.8 is also observed in single bubble growth.
-Scriven’s result is valid for a spherically symmetric concentration field, bubble radius much larger than the initial radius, constant solubility, and, far away from the bubble, constant concentration and zero velocity.
-This solution is confirmed in cases where the thickness $\delta_{\mathrm{H_{2}}}$ of the diffusion boundary layer, surrounding the bubble is small compared with the diameter of the bubbles.
-Given the constraints of the analytical solution, the discrepancies in our simulation can be explained.
-The hydrogen flux at the boundary determined by current density influences the growth exponent increasing the averaged hydrogen concentration over the electrode wall.
-At the beginning stage, the growth exponent is not close to $1/2$.
+The inertia-driven and diffusion-driven growth regimes correspond to slopes of 1 and $1/2$, respectively.
+As shown in Fig.˜6, the transition to diffusion-driven growth is evident, but the expected inertia-driven regime is missing.
+The steepest slope is observed only during the initial growth stage at the highest current density, and is close to $0.8$.
+The diffusion-driven growth slope of $1/2$ is assumed to be valid for a spherically symmetric concentration field, a bubble radius much larger than the initial radius, constant solubility, and, far from the bubble, constant concentration and zero velocity.
+This solution is confirmed in cases where the thickness $\delta_{\mathrm{H_{2}}}$ of the diffusion boundary layer surrounding the bubble is small compared with the diameter of the bubbles.
+Given these constraints of the analytical solution, the discrepancies in our simulation can be explained.
+The hydrogen flux at the boundary, which is determined by current density, influences the growth exponent by increasing the averaged hydrogen concentration on the electrode wall.
+At the beginning stage, the growth exponent is larger than $1/2$.
 The bubble radius is relatively small compared to the thickness of the diffusion boundary layer.
-For high current densities $I=1000\mathrm{A/m^{2}}$, the enhanced hydrogen flux elevates the supersaturation level and creates a strong diffusion boundary layer.
-The diffusion layer continuously provides $\mathrm{H_{2}}$ for bubble growth, which by consequence leads a much faster $\mathrm{H_{2}}$ production than consumption and a growth exponent exceeding $1/2$.
-Conversely, lower current densities reduce hydrogen availability, can not fulfill the consumption of bubble growth, leading to sub $1/2$ exponents.
+At high current densities, such as $I=1000\,\mathrm{A/m^{2}}$, the increased hydrogen flux raises the supersaturation level and generates a strong diffusion boundary layer.
+This layer continuously supplies $\mathrm{H_{2}}$ for bubble growth, resulting in a production rate that surpasses consumption and yields a growth exponent greater than $1/2$.
+Numerically, the stronger flux produces a steeper $\mathrm{H_{2}}$ concentration gradient, which enhances mass transfer across the interface.
+In contrast, lower current densities limit hydrogen availability, making it insufficient to sustain bubble growth and leading to sub-$1/2$ exponents.
+The decreasing trend of the growth exponent with lower current densities confirms this behavior.
 After the initial growth stage, as the bubble radius increases, the $\delta_{\mathrm{H_{2}}}$ is much smaller compared with the bubble radius.
-The growth exponents for different boundary condition are all approaching $1/2$, which indicates that the overall growth is diffusion controlled.
-Figure 5: The dissolved hydrogen concentration for a bubble growing at different times (The current density is $I=\mathrm{1000A/m^{2}}$).The color lengend represents the numerical $\mathrm{H_{2}}$ concentration.
-Figure 6: Bubble radius growth for bubbles with $\theta=90\degree$ for different current densities. The units of $I$ are $\mathrm{A/m^{2}}$
-Figure 7: The growth bubble radius slope and growth exponent variation.
-To further compare the numerical results with experimental data, additional simulations are performed with a smaller contact angle ($\theta=35^{\circ}$), which is the same value from the experimental study(see Glas and Westwater (1964)).
-From the analytical solution $R=2\beta t^{1/2}$, the growth rate $\beta$ could be deduced.
-The results in terms of the influence of current density on the growth rate are reported in fig.
-8.
-The first comparison indicates that the growth rate increases with rising current density, attributed to the high concentration of dissolved hydrogen near the electrode.
-This observation is supported by both the numerical simulations conducted by Gennari et al.
-(2022) and experimental results from Glas and Westwater (1964).
-The observation can be explained by the current density controlling the hydrogen flux at the electrode, which drives bubble growth.
-Figure 8: Comparison of growth rate between experimental and numerical results.
-We find that a bubble with a contact angle of $\theta=90^{\circ}$ grows more rapidly than one with a contact angle of $\theta=35^{\circ}$.
-This difference can be explained by considering the geometric difference of different contact angles.
-The mass transfer through diffusion occurs only through the interface between the gas and liquid phase.
-In the $\theta=90^{\circ}$ scenario, the bubble elongates more in the r-direction (parallel to the electrode wall), thereby increasing its exposure to the high-concentration region close to the electrode.
-So, the mass transfer for the bubble with $\theta=90\degree$ would be higher.
-A qualitative comparison between $\theta=90\degree$ and $\theta=35\degree$ is shown in fig.
-9 and fig.
-10.
-It is notable that the simulation of a single bubble overpredicts the growth rate compared to the experimental measurements.
-This discrepancy is observed not only in the present simulation but also in the results from Gennari et al.
-(2022).
-In fact, there are at least four active nucleation sites on the electrode surface from the experiment of Glas and Westwater (1964).
-This suggests that experimental observations of single bubble growth are generally influenced by neighboring bubbles.
-Depending on the material and surface properties of the electrode, if multiple nucleation sites are active, the dissolved hydrogen in the liquid is consumed simultaneously by all of them.
-Consequently, the relative growth rates are smaller than in the case of a single nucleation site, where the entire amount of $\mathrm{H_{2}}$ produced by the electrode is available for a single bubble on the wall.
-This mechanism explains the overprediction in the simulations.
-To further investigate the effect of variations in nucleation site distribution, a 3D simulation would be necessary, as the scenario would no longer be axisymmetric.
-Figure 9: Bubble shape and contour of $\mathrm{H_{2}}$ concentration for $\theta=90\degree$.
-Figure 10: Bubble shape and contour of $\mathrm{H_{2}}$ concentration for $\theta=35\degree$.
-3.2 3D simulation extension
-To extend the numerical scheme to 3D simulation, we set the domain as a cubic (fig.
-3).
+The growth exponents for different boundary conditions are all approaching $1/2$, which indicates that the overall growth is diffusion-controlled.
+Figure 5: Frames show the dissolved hydrogen concentration for a bubble growing at the contact angle of $\theta=90\degree$(with the current density of $I=\mathrm{1000A/m^{2}}$).
+The color legend represents $\mathrm{H_{2}}$ concentration, while the upper limit is $0.2\mathrm{mol/m^{3}}$.
+Figure 6: The bubble radius growth slope and growth exponent variation.
+3.2 Three-dimensional simulation
+To extend the numerical scheme to three-dimensional simulation, we set the domain as a cube (Fig.˜3).
 And the boundary condition would be modified to suit the domain.
-The electrode will be placed in the middle of the domain bottom(XY plane), which will give the same boundary condition of hydrogen flux as axisymmetric simulation, see eq.
-4.
-The top plane(XY) will be set as an outflow boundary condition, and the surrounding area (XZ and YZ planes) adopt a no-slip wall boundary condition.
-In this section, a single bubble is modeled in 3D to validate the 3D applicability of the numerical method.
-We compare the bubble growth prediction of the 3D simulation to that of the corresponding axisymmetric simulation.
-Simulations of configurations No.1 and No.7 (Table 1) with three different current densities are listed for comparison.
-The contact angle is $90\degree$.
-Same bubble growth slope from section 3.1 are observed in fig.
-11.
-Since the convergence has been verified in section 3.1, the good agreement with the axisymmetric simulation results proves the compatibility of 3D simulation.
-Figure 11: Bubble radius growth for bubbles with $\theta=90\degree$ for different current densities. The units of $I$ are $\mathrm{A/m^{2}}$.
+The electrode will be placed in the middle of the domain bottom boundary ($xy$ plane), which will give the same boundary condition of hydrogen flux as an axisymmetric simulation, see Eq.˜4.
+The top $xy$ plane will be set as an outflow boundary condition, and on the lateral boundaries ($xz$ and $yz$ planes), a no-slip boundary condition is set.
+A single bubble is modeled in three dimensions to validate the applicability of the numerical method.
+We compare the bubble growth prediction of the three-dimensional simulation to that of the corresponding axisymmetric simulation.
+Specifically, simulations of configurations No.
+1 and No.
+6 (Table˜1) are compared.
+The results show that the bubble growth slopes obtained from the axisymmetric simulations are identical to those from the three-dimensional simulation.
+Since the numerical convergence has been verified in Section˜3.1, this good agreement confirms the consistency of the three-dimensional simulation.
+Consequently, the three-dimensional configuration is used for the subsequent simulations.
 4 Result and discussion
+4.1 Single bubble growth
+Based on the experiments of Glas1964, both the current density and the electrode surface covering (determined by the contact angle $\theta$) influence the bubble growth.
+A notable observation is that the contact angle is variable.
+Glas1964 found that the change is large, from about 70°to 20°.
+However, for simplicity, our numerical model assumes a constant contact angle throughout the bubble growth evolution.
+For further comparison, another series of simulations is performed with a smaller contact angle ($\theta=35^{\circ}$), which corresponds to a hydrophilic electrode surface.
+From the analytical solution $R=2\beta t^{1/2}$, the growth rate $\beta$ can be determined.
+The fitting method used by Glas1964 involves plotting the bubble radius $R$ versus $t^{1/2}$, from which $\beta$ is computed directly from the slope.
+Using the same historical fitting approach, the numerical results showing the influence of current density on the growth rate are presented in Fig.˜7.
+Because the bubble growth exponent does not perfectly match $1/2$ during the early growth stage (as discussed in Fig.˜6), we performed three local fittings at different times for each case, with the resulting error bar indicating the variability among these estimates.
+The results clearly show that the growth rate increases with increasing current density.
+This trend is expected, as the current density directly dictates the hydrogen flux at the electrode surface, which drives bubble growth through higher concentrations of dissolved hydrogen near the electrode.
+Figure 7: Comparison of growth rate of a single bubble between experimental and numerical results.
+Then, we find that a bubble with a contact angle of $\theta=90^{\circ}$ grows faster than one with a contact angle of $\theta=35^{\circ}$.
+This difference can be explained by considering the influence of the contact angle on the bubble shape and placement.
+Mass transfer through diffusion occurs only through the interface between the gas and liquid phases.
+In the $\theta=90^{\circ}$ scenario, the bubble elongates more along the electrode wall, thereby increasing its exposure to the high-concentration region close to the electrode.
+As a result, the effective diffusive mass transfer to the bubble with $\theta=90\degree$ is enhanced.
+Furthermore, Fig.˜8 presents the growth rate trend deduced from several different contact angles.
+The simulations cover the range from $35^{\circ}\leq\theta\leq 90^{\circ}$ for the same initial nucleation diameter.
+This range is relevant as most electrode surfaces are designed to be hydrophilic to facilitate the removal of attached bubbles.
+The results indicate that the bubble-growth coefficient generally increases with the contact angle, until the upper limit of our simulations of $\theta=90^{\circ}$.
+However, as the contact angle approaches the lower limit, the bubble-growth coefficient is no longer dominated by the contact angle.
+In some instances, such as at a current density of $I=500\,\mathrm{A/m^{2}}$, this trend is even reversed.
+While we have no explanation for this reversal, we attribute the weakening influence of the contact angle near the lower limit to a geometrical effect.
+Indeed, as the contact angle decreases, the bubble geometry changes only a little while it approaches a perfect sphere, and the influence of the contact angle on growth becomes less significant.
+Consequently, simulation results for contact angles smaller than $\theta=35^{\circ}$ are not shown here.
+Figure 8: Comparison of growth rate of a single bubble between experimental and numerical results for various contact angles.
+A significant discrepancy in the effect of current density is also noted when comparing the simulated and experimentally measured bubble-growth coefficients $\beta$, as shown in Fig.˜7.
+For the $\theta=35^{\circ}$ case, the numerical model underestimates $\beta$ at low current densities.
+The deviation is more complex for $\theta=90^{\circ}$, where $\beta$ is underestimated at low current densities but is conversely overestimated at higher current densities.
+This discrepancy is observed not only in the present simulation but also in the results of Gennari2022.
+The underestimation is likely due to the difference between the numerical setting and the experimental conditions.
+Glas1964 established a nonlinear relationship between the growth coefficient $\beta$ and the current density $I$, as the driving force (local supersaturation) inherently depends on the current density.
+This model $\beta=I^{a},a<1$ effectively describes the experimental minor change of $\beta$ observed for $(100<I<500\,\mathrm{A/m^{2}})$.
+In contrast, our numerical work assumes that the rate of $\mathrm{H_{2}}$ production is directly proportional to the electrical current.
+This simplified setting cannot perfectly replicate the real situation, where bubble production at any site is intermittent in practice.
+Furthermore, the simulation uses a typical nucleation time of $0.02\mathrm{s}$.
+This time point determines the concentration layer of dissolved hydrogen when bubble growth begins.
+However, the nucleation time in experiments can vary between $0.02\mathrm{s}$ to $0.1\mathrm{s}$ (Glas1964).
+The difference in nucleation time at low current densities results in a weaker hydrogen concentration boundary layer than the experimental boundary layer, and therefore a smaller growth coefficient.
+To explain the overestimation observed in the simulations, two major differences between the simulated and experimental conditions should be noted.
+First, the contact angle is found to change from $70\degree$ to $20\degree$ in the experiment.
+Since our simulations utilize a constant contact angle, a perfect match across all experimental conditions is not expected.
+However, it is noteworthy that the discrepancy is minimal when comparing the simulation results within the experimentally observed range of $35\degree\leq\theta\leq 70\degree$.
+Second, obtaining a truly isolated, single active nucleation site on an electrode surface is challenging in both modern experiments and earlier studies (Glas1964).
+This suggests that experimental measurements of presumed single-bubble growth may, in fact, be influenced by neighboring bubbles, such as a bubble carpet or multiple nucleation sites.
+When multiple nucleation sites are active (a condition dependent on the electrode material and its surface properties), the dissolved hydrogen in the liquid is consumed simultaneously by several bubbles.
+Consequently, the individual growth rates are reduced compared with the idealized case of a single nucleation site, which benefits from the full hydrogen flux from the electrode, contributing to the growth of one bubble alone.
 4.2 Multiple bubbles growth
-There is a notable decrease in the growth rate and exponent of hydrogen bubbles when multiple nucleation sites are active from the experiment of Glas and Westwater (1964).
-To reproduce this key observation, this section numerically investigates the mutual bubble interactions for multiple nucleation sites.
-Instead of a single bubble, we model 2 bubbles and 4 bubbles growing at the electrode.
-To highlight the effect of nucleation sites, a higher boundary flux is applied, with the current density set to $\mathrm{1000A/m^{2}}$.
-The simulation results for the average bubble radius growth of two and four bubbles are presented in fig.
-12.
-The growth of multiple nucleation sites is mutually suppressed in the stage before bubble coalescence.
+There is a notable decrease in the growth rate and exponent of hydrogen bubbles when multiple nucleation sites are active (Glas1964).
+This Section numerically investigates the mutual bubble interactions for multiple nucleation sites.
+Instead of a single bubble, we model two bubbles and four bubbles growing at the same electrode.
+Since the influence of the contact angle on bubble growth has already been discussed in Section˜4.1, the results presented in this section focus on bubbles growing at a contact angle of $\theta=90\degree$.
+The growth of multiple nucleation sites is mutually suppressed in the stage before bubble coalescence from the growth slope Fig.˜9.
 During this stage, hydrogen is consumed by all the bubbles simultaneously, leading to a reduced growth rate and exponent.
-The more active nucleation sites, the slower the growth, as shown in the comparison between experimental and simulation results in fig.
-13.
-However, an unexpected behavior of the growth seems puzzling.
-Intuitively, merged bubbles should approximate the growth dynamics of a single nucleation site, leading to convergent growth across all cases.
-Contrary to this assumption, even after bubble coalescence($t>t_{m}$), differences in the growth persist.
-Figure 12: Bubble growth of configurations No.7 to No.9. The current density is $I=\mathrm{1000A/m^{2}}$ and the slope represents the growth exponent.
-Figure 13: Influence of nucleation sites on the growth rate.
+The larger the number of active nucleation sites, the slower the growth.
+However, an unexpected behavior of the bubbles’ growth seems puzzling.
+Intuitively, once the bubbles merge, they should approximate the growth dynamics of a single bubble case.
+Contrary to this assumption, even after bubble coalescence($t>t_{m}$), differences in the growth persist for a period.
+Then the growth curves of all cases converge, as shown in Fig.˜9.
 A possible explanation lies in the relative position of bubbles within the hydrogen concentration layer.
-As shown in fig.
-12, the trend remains consistent: the more active nucleation sites, the slower the growth, regardless of the stage (before or after bubble coalescence).
-(a) Single nucleation site at merge time $t_{m}$.
-(b) Four nucleation sites at merge time $t_{m}$.
-Figure 14: Comparison of bubble growth between different nucleation sites.
-The current density is 1000 A/m2.
-The view direction is the same as in fig.
-3.
-The hydrogen contour on the intersecting plane (YZ) in the middle of the bubble is transferred to the back plane.
+As shown in Fig.˜9, the trend remains consistent: the more active nucleation sites, the slower the growth, regardless of the stage (before or after bubble coalescence).
+Figure 9: Growth of multi-bubbles at the contact angle of $\theta=90\degree$. The current density is $\mathrm{1000A/m^{2}}$ and the slope represents the growth exponent.
 The key reason for the different growth behavior after coalescence is the bubble radius.
@@
 This is because the hydrogen flux at the boundary is sufficiently high, making the amount of hydrogen consumed by bubble growth negligible.
-Then, the crucial factor is immersion depth: bigger bubbles immersed shallower into the hydrogen concentration layer, leading to a lower mass transfer rate, which means a smaller growth exponent, as illustrated in fig.
-12.
-Supporting evidence emerges from direct comparisons: at the same time $t_{m}$, the four-nucleation-site configuration produces marginally larger bubbles than the single-site case, see fig.
-14.
-However, the constant hydrogen flux induces a countervailing effect as illustrated in the snapshots fig.
-14, the bubble surface is colored by the mass transfer rate.
-A higher mass transfer rate slows the increase of the hydrogen concentration boundary layer for the four-nucleation-site case, enabling shallower bubble immersion after merge.
-This feedback mechanism drives a gradual decrease in growth exponent, ultimately causing convergence across nucleation-site configurations.
-As shown in fig.
-12, the growth exponent gradually approaches a slope of 0.5.
-To quantitatively assess mass transfer, we analyze the time-dependent bubble Sherwood number defined in eq.
-8, which quantifies mass transport efficiency to the bubble.
-The initial condition $\zeta=1$, delay the plot until $t>t_{n}$, which is shown in fig.
-15.
+Then, the crucial factor is immersion depth: bigger bubbles immersed less deeply into the hydrogen concentration layer, leading to a lower mass transfer rate, which means a smaller growth exponent, as illustrated in Fig.˜9.
+Supporting evidence emerges from direct comparisons: at the same time $t_{m}$, the four-nucleation-site configuration produces marginally larger bubbles than the single-site case, as shown in Fig.˜10.
+However, the constant hydrogen flux induces a countervailing effect.
+As illustrated by the snapshots in Fig.˜10, where the bubble surface is colored by the local mass transfer rate, a higher overall mass transfer rate slows the increase of the hydrogen concentration boundary layer for the four-nucleation-site case.
+This effect enables shallower bubble immersion after coalescence.
+This feedback mechanism drives a gradual decrease in the growth exponent, ultimately causing convergence across different nucleation-site configurations.
+As shown in Fig.˜9, the growth exponent gradually approaches a slope of $1/2$.
+(a) Single nucleation site at merge time $t_{m}$.
+(b) Four nucleation sites at merge time $t_{m}$.
+Figure 10: Frames show the comparison between the single bubble and four bubbles at the contact angle of $\theta=90\degree$.
+The current density is 1000 A/m2.
+The view direction is the same as in Fig.˜3.
+The hydrogen contour on the intersecting plane (yz) passing through the middle of the bubble is projected to the back plane.
+The bubble interface is colored according to the local mass-transfer rate.
+To quantitatively assess mass transfer, Fig.˜11 allows to analyze the time-dependent bubble Sherwood number defined in Eq.˜8, which quantifies mass transport efficiency to the bubble.
 It should be noticed that multiple nucleation sites exhibit elevated $\mathrm{Sh_{b}}$ initially.
-A decrease occurs after bubble merge($t_{m}$), which causes the $\mathrm{Sh_{b}}$ to be smaller than the one for the single nucleation site.
-The plot of $\mathrm{Sh_{b}}$ cross-verified the explanation of growth exponent above.
+A decrease occurs after bubble coalescence($t_{m}$), which causes the $\mathrm{Sh_{b}}$ to be smaller than the one for the single nucleation site.
+The plot of $\mathrm{Sh_{b}}$ cross-verified the explanation of the difference in growth rate above.
 First, the mutual suppression.
@@
 Second, the immersion depth: lower after-merge $\mathrm{Sh_{b}}$ directly correlates with diminished mass transfer rates for multiple nucleation sites, consistent with the immersion depth arguments.
-Figure 15: Corresponding Sherwood number for simulations of configurations No.7 to No.9.
+Figure 11: The Sherwood number for simulations with different numbers of nucleation sites. The current density is $1000\mathrm{A/m^{2}}$.
+As shown in the comparison between experimental and simulation results in Fig.˜12, the growth rates obtained from the multiple-bubble cases exhibit better agreement with experimental data than those from the single-bubble case (Fig.˜7).
+This observation supports the earlier conjecture regarding the influence of neighboring bubbles.
+Although the reported experimental growth rate used for comparison was acquired for a single bubble attached to the electrode, likely, small bubbles forming underneath were not captured during measurement.
+Consequently, the growth rates predicted by the multiple-bubble simulations may more accurately represent the actual experimental conditions.
+Furthermore, the experimental data itself exhibits a considerable standard deviation, and the growth coefficient $\beta$ was obtained through a fitting process without associated error bars.
+This experimental uncertainty presents an alternative explanation for the observed differences between simulation and experimental measurement.
+Figure 12: Comparison of growth rate for multiple nucleation sites with the contact angle of $\theta=90\degree$.
+Figure 13: Comparison of growth rate for single site and multiple nucleation sites with the contact angle of $\theta=90\degree$.
+To comprehensively assess the influence of nucleation sites, Fig.˜13 compares the average bubble growth coefficients $\beta$ for one, two, and four bubbles.
+In addition to the expected suppression of the growth coefficient, the results reveal that the relative impact of multiple nucleation sites becomes more significant at higher current densities.
+This observation may help explain the earlier discrepancies between simulations and experiments: the higher current density amplifies the effect of other parameters, such as contact angle and the density of nucleation sites.
+Notably, the overestimation of the bubble-growth coefficient is remarkable at the current density of $I=1000\,\mathrm{A/m^{2}}$, whereas at lower current densities $(300<I<800\,\mathrm{A/m^{2}})$, the numerical predictions are comparable to the experimental measurements.
 4.3 Bubble detachment
-The removal of growing bubbles is crucial for enhancing the efficiency of the electrolysis process (Angulo et al., 2020).
-The bubble detachment radius is approximately equal to the Fritz radius (Stephan and Vogt, 1979)
-$R_{det}=0.6\theta\sqrt{\frac{\sigma}{(\rho_{c}-\rho_{d})g}},$ (37)
-where $\theta$ the contact angle in radians.
-The general contact angle for the bubble at the electrode would be smaller than $\pi/2$.
+The removal of attached bubbles is crucial for enhancing the efficiency of the electrolysis process (Angulo2020).
+In the absence of contact angle hysteresis, the bubble tends to spread on the electrode surface with constant speed.
+The force balance determines the relation between the volume-equivalent detachment radius (Fritz radius) and the value of the contact angle $\theta$ giving(Stephan1979),
+$R_{det}=0.6\,\theta\sqrt{\frac{\sigma}{(\rho_{c}-\rho_{d})g}}.$ (37)
 This analytical solution is derived by balancing the buoyancy force for a perfect sphere-shaped bubble $F_{b}=(4/3)\pi R^{3}(\rho_{c}-\rho_{d})g$ with the capillary force for a bubble attached to the electrode.
-The contact angle controls the magnitude of surface tension effects and varies in connection with contact line dynamics, which are still, in many circumstances, an object of investigation.
-Molecular dynamics simulations with either Lennard-Jones intermolecular potentials or more realistic water molecule models give very different results, see Lācis et al.
-(2022).
-Therefore, in the present study, the theoretical value is adopted as a reference to estimate the order of magnitude rather than for precise comparison.
-To begin with, a simulation of single-bubble detachment is shown in fig.
-16.
-The numerical configuration refers to the No.1 and No.4 in Table 1.
-As observed, once the mass transfer is considered (at $t=20\mathrm{ms}$), the bubble starts to grow.
+To begin with, a simulation of single-bubble detachment is performed.
 Initially, the bubble remains nearly spherical due to its small size, which results in a negligible buoyancy force.
-However, as the bubble radius increases, deformation becomes evident, particularly after $t=52.1\mathrm{ms}$.
-Eventually, buoyancy overcomes surface tension, leading to detachment from the electrode surface.
-Figure 16: Frames show the detachment evolution($\theta=90^{\circ}$).
-As previously mentioned, eq.
-37 establishes a relationship between the detachment radius and the contact angle.
-To further investigate this dependency, another case with a contact angle $\theta=35\degree$ was also simulated, see fig.
-17.
-Compared to the case $\theta=90\degree$, the detachment time is much shorter, and the detachment radius is much smaller.
-fig.
-18 presents a good agreement between the numerical simulation and the Fritz radius.
-Figure 17: Frames show the detachment evolution($\theta=35^{\circ}$).
-Figure 18: Bubble detachment radius versus contact angle for a single bubble
-Additionally, it has been confirmed that mutual interactions between bubbles influence the detachment process.
-Bashkatov et al.
-(2024) reported that bubble coalescence in the presence of dual bubbles can lead to significantly earlier detachment and a smaller detachment radius compared to cases where only buoyancy effects are considered.
-To validate this finding, we perform a series of simulations involving bubble detachments for multiple nucleation sites.
-The results of 3D simulations with four nucleation sites are presented in fig.
-19.
-Furthermore, varying the number of nucleation sites demonstrates that bubble coalescence indeed accelerates detachment, as shown in fig.
-20.
-As a result, the bubble detachment radius is also smaller, see fig.
-21.
-A plausible mechanism behind this acceleration could be the following.
+However, as the bubble radius increases, deformation becomes evident.
+Eventually, buoyancy overcomes surface tension, leading to the detachment from the electrode surface.
+Figure 14: Bubble detachment radius versus contact angle for a single bubble(with the current density of $I=\mathrm{1000A/m^{2}}$).
+Figure 15: Frames show the detachment evolution for a single bubble at different contact angles(with the current density of $I=\mathrm{1000A/m^{2}}$).
+The corresponding detachment times ($t_{d}$) are $104.2\mathrm{ms}$ for $\theta=35^{\circ}$, and $243.2\mathrm{ms}$ for $\theta=90^{\circ}$.
+The hydrogen contour on the intersecting plane (yz) passing through the middle of the bubble is projected to the back plane.
+The bubble interface is colored according to the local mass-transfer rate.
+As previously mentioned, Eq.˜37 establishes a linear relationship between $R_{d}$ and $\theta$.
+To further investigate this dependency, a series of cases with varying contact angle from $\theta=35\degree$ to $\theta=90\degree$ are simulated, see Fig.˜14.
+As expected, the numerical results also exhibit a linear relationship.
+The Fritz radius is computed using the same reduced surface tension.
+However, the simulation predicts larger detachment radii compared to the theoretical value.
+The pinning effect of contact angle hysteresis can explain this deviation.
+For detachment to occur, the contact line must recede (shrink inwards).
+However, contact angle hysteresis pins the contact line.
+The bubble cannot simply shrink at its base; instead, its neck must stretch and thin until the surface tension force can no longer hold it.
+Since the theoretical Fritz radius is derived from a simplified force balance that neglects contact angle hysteresis, the theoretical value is expected to be smaller than the detachment radius predicted by the simulation.
+To demonstrate the effect of contact angle on detachment radius, we take the most contrasting cases, $\theta=90^{\circ}$ (hydrophobic) and $\theta=35^{\circ}$ (hydrophilic), as an example.
+As shown in Fig.˜15, the bubble with the smaller contact angle ($\theta=35^{\circ}$) detaches significantly faster and at a much smaller radius.
+Additionally, it has been shown that mutual interactions between bubbles influence the detachment process.
+Bashkatov2024 observed that bubble coalescence in the presence of dual bubbles can lead to significantly earlier detachment and a smaller detachment radius compared to cases where only buoyancy effects are considered.
+As illustrated in Fig.˜16, the presence of multiple bubbles indeed accelerates the detachment.
+Consequently, the detachment radius is reduced, as demonstrated in Fig.˜17.
+For a direct visual comparison, representative frames from simulations with multiple bubbles at contact angles of $\theta=90^{\circ}$ and $\theta=35^{\circ}$ are provided in Figs.˜18 and 19.
+Figure 16: Bubble detachment times for different nucleation sites and contact angles (with the current density of $I=\mathrm{1000A/m^{2}}$).
+Figure 17: Bubble detachment radius for different nucleation sites and contact angles (with the current density of $I=\mathrm{1000A/m^{2}}$).
+Figure 18: Frames show the detachment evolution for multi-bubbles at the contact angle of $\theta=90\degree$ (with the current density of $I=\mathrm{1000A/m^{2}}$).
+The corresponding detachment times ($t_{d}$) are $243.2\mathrm{ms}$ for 1 bubble, $241.4\mathrm{ms}$ for the 2-bubbles case, and $235.5\mathrm{ms}$ for the 4-bubbles case.
+The hydrogen contour on the intersecting plane (yz) passing through the middle of the bubble is projected to the back plane.
+The bubble interface is colored according to the local mass-transfer rate.
+Figure 19: Frames show the detachment evolution for multi-bubbles at the contact angle of $\theta=35\degree$ (with the current density of $I=\mathrm{1000A/m^{2}}$).
+The corresponding detachment times ($t_{d}$) are $104.2\mathrm{ms}$ for 1 bubble, $94.6\mathrm{ms}$ for the 2-bubbles case, and $67.6\mathrm{ms}$ for the 4-bubbles case.
+The hydrogen contour on the intersecting plane (yz) passing through the middle of the bubble is projected to the back plane.
+The bubble interface is colored according to the local mass-transfer rate.
+A plausible mechanism behind this acceleration is as follows.
 The coalescence-induced shape perturbations generate an initial momentum favoring premature detachment from the electrode surface.
-It is observed in both our simulation and Bashkatov’s experiment that the coalescence of bubbles results in an initial jump-off of the merged bubble.
-This can be attributed to the released surface energy during coalescence (Zhang et al., 2024).
+It is observed in Bashkatov’s experiment that the coalescence of bubbles results in an initial jump-off of the merged bubble.
+This can be attributed to the released surface energy during coalescence (Zhang2024).
 The released energy is partly dissipated by the bubble oscillations, working against viscous drag.
-When in the proximity to the surface, the remaining energy is converted to kinetic energy, driving the resultant (merged) bubble to jump off the electrode (Lv et al., 2021).
-This kinematic energy can cause bubble departure at smaller radii than in the buoyancy-driven scenario, as observed in our simulation results.
-However, the capillary wave due to energy dissipation is not observed in our simulation because of the mesh resolution constraint.
+When in the proximity to the surface, the remaining energy is converted to kinetic energy, driving the resultant (merged) bubble to jump off the electrode (lv2021).
+This kinetic energy can cause bubble departure at smaller radii than in the purely buoyancy-driven scenario, as observed in our simulation results.
+However, the capillary waves due to energy dissipation are not observed in our simulations because of the mesh resolution constraint.
 Further numerical investigations are needed to distinguish the respective roles of buoyancy and interfacial effects in this process.
-Figure 19: snapshots of the detachment evolution for 4 nucleation sites (Configuration No.9).
-Figure 20: Bubble detachment times for different nucleation sites and contact angles.
-Figure 21: Bubble detachment radius for different nucleation sites and contact angles.
 5 Conclusion
@@
 The following conclusions are drawn:
-Bubble growth dynamics The bubble contact angle significantly influences growth behavior by dictating the geometric position of the bubble within the evolving hydrogen concentration profile.
-This directly modulates local mass transport rates.
-For multi-bubble systems, mutual suppression of mass transport occurs due to simultaneous hydrogen consumption by neighboring bubbles, as quantified by reductions in Sherwood number (fig.
-15).
-While simulations overpredict growth rates compared to experimental observations (fig.
-13), they adhere closely to the scaling law of Scriven’s solution ($R\propto t^{1/2}$).
-Bubble detachment and coalescence behavior Detachment radii align with Fritz’s formula, demonstrating a decrease for smaller contact angles.
-Multi-bubble systems exhibit earlier detachment with reduced radii compared to isolated bubbles.
-t is worth noting that coalescence reduces the detachment time, thus potentially improving the mass transfer efficiency of the electrochemical system.
-Prospective The present work focuses on single growth-detachment cycles due to computational constraints.
-Tracking deformable interfaces remains resource-intensive, limiting simulations to short time spans ($t\sim 0.2$s).
-While comparable to prior studies using rigid bubble models (Sepahi et al., 2022), broader parametric investigations (e.g., current density variations, contact angle distributions) are needed to strengthen statistical conclusions.
-Future efforts are needed to develop a new scheme with subgrid model, which will help to enable long-term simulations.
-Additionally, it should be noted that the experimental data used for comparison are not fully aligned with the study’s objectives.
-We will design and implement a more sophisticated experimental setup that will provide observational proof of our work (e.g., continuous single bubble detachment), thus increasing the reliability and impact of the 3D numerical study.
+Bubble growth dynamics
+The bubble contact angle significantly influences growth behavior by dictating the geometric position of the bubble within the evolving hydrogen concentration profile.
+This geometric position directly modulates the local mass-transfer rates.
+For multi-bubble systems, mutual suppression of mass transport occurs due to the simultaneous hydrogen consumption by neighboring bubbles, a phenomenon quantified by reductions in the Sherwood number ($\mathrm{Sh}_{b}$), as shown in Fig.˜11.
+Specifically, increasing the number of nucleation sites leads to a smaller individual bubble Sherwood number.
+While some numerical overprediction and underprediction of growth rates exist compared to experimental observations (Fig.˜12), the simulations adhere closely to the scaling law of Scriven’s solution ($R\propto t^{1/2}$).
+Bubble detachment and coalescence behavior
+Detachment radii align with Fritz’s formula, demonstrating a linear relation between the volume-equivalent radii and the contact angles.
+Multi-bubble cases exhibit earlier detachment with reduced radii compared to isolated bubbles.
+It is worth noting that coalescence reduces the detachment time, thus potentially improving the mass transfer efficiency of the electrochemical system.
+Outlook
+The present work focuses on single growth-detachment cycles due to computational constraints.
+Tracking deformable interfaces remains resource-intensive, limiting simulations to short time spans ($t\sim 0.2\,\mathrm{s}$).
+While comparable to prior studies using rigid bubble models (sepahi2022a), broader parametric investigations (e.g., current density variations, contact angle distributions) are necessary to strengthen statistical conclusions.
+For instance, a statistically steady Sherwood number at the electrode $\mathrm{Sh_{e}}$ has not been investigated in this work.
+Future efforts are required to develop a new scheme with a larger stable time step, which will help to enable long-term simulations.
 Declaration of competing interest